### v2

# From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?

Abstract.
Engagement with misinformation on social media poses unprecedented threats to societal well-being, particularly during health crises when susceptibility to misinformation is heightened in a multi-topic context. This paper focuses on the COVID-19 pandemic and addresses a critical gap in understanding online engagement with multi-topic misinformation at two user levels: news sharers who share source news items on social media and post viewers who engage with online news posts. To this end, we conduct a comprehensive analysis of fact-checked source news claims related to COVID-19 and their associated posts on X, through the lens of topic diversity and conspiracy theories. We find that false news, particularly when accompanied by conspiracy theories, exhibits higher topic diversity than true news. At the news sharer level, false news has a longer lifetime and receives more posts on X than true news. Additionally, the integration of conspiracy theories is significantly associated with a longer lifetime for COVID-19 misinformation. However, topic diversity has no significant association with news sharer engagement in terms of news lifetime and the number of posts. At the post viewer level, contrary to the news sharer level, news posts characterized by heightened topic diversity receive more reposts, likes, and replies. Notably, post viewers tend to engage more with misinformation containing conspiracy narratives: false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. Our findings suggest that news sharers and post viewers exhibit different engagement patterns on social media regarding topic diversity and conspiracy theories, offering valuable insights into designing targeted misinformation intervention strategies at both user levels.
1. Introduction
The spread of misinformation on social media platforms has emerged as one of the most significant global risks in recent years, posing far-reaching consequences for societies and individuals (Ecker et al., 2022). Health misinformation, compared to misinformation in other domains, carries unique dangers due to its potential to cause harm to human bodies and lives (Chen et al., 2022). This issue is particularly prevalent during public health crises, such as the COVID-19 pandemic, when heightened emotions like fear and anxiety, uncertainty, and increased information consumption create fertile ground for the dissemination of false or misleading information (Chuai and Zhao, 2022; Solovev and Pröllochs, 2022; Chuai et al., 2022; Chen et al., 2023; Freiling et al., 2021). Throughout the COVID-19 pandemic, a substantial amount of health misinformation regarding the virus, its origin, and potential treatments or prevention strategies spread widely on social media (Erokhin et al., 2022; Sharma et al., 2022; Teng et al., 2022). Especially, the widespread COVID-19 vaccine misinformation on social media increases beliefs in the ineffectiveness and side effects of vaccinations, contributing to individuals’ vaccination hesitancy and refusal, thereby increasing the risk of infection (Sharma et al., 2022; Pierri et al., 2022; Enders et al., 2022).
Furthermore, during health crises, various aspects of society and individuals become interconnected. The discussions traverse multiple domains and demonstrate strong interconnectedness, contributing to a rich and diverse informational landscape (Miani et al., 2022; Wang et al., 2022). However, there remains a gap in understanding online engagement with misinformation within this multi-topic context during health crises. A comprehensive analysis of 13.9 million COVID-19 posts on X reveals discussions spanning more than 20 topics across the societies, encompassing areas such as politics, economy, and racism (Chandrasekaran et al., 2020). The diversity of topics within information content (i. e., multi-topic information) plays an important role in shaping perceptions and assessments of credibility through the mechanism of associative inference (Lee et al., 2023). Associative inference is an adaptive process that enables individuals to link related information and form novel connections, even in the absence of direct experience, thereby increasing individuals’ susceptibility to online misinformation (Lee et al., 2023). Moreover, the associative inference around online misinformation can evolve within a reinforcing cycle. Individuals collaborate in constructing a misleading version of reality, leading to offline actions that further reinforce this manufactured reality (Prochaska et al., 2023). The interconnected and associative multi-topic discussions during health crises underscore the importance of exploring the role of topic diversity in engagement with misinformation on social media.
Additionally, conspiracy theories (CTs) often emerge during health crises, attempting to explain the causes of significant public events as secret plots by powerful and malicious groups (Douglas and Sutton, 2018; Kou et al., 2017; Suarez-Lledo and Alvarez-Galvez, 2021). The integration of conspiracy theories within misinformation narratives serves as an effective strategy to deliberately interconnect various topics (Miani et al., 2022; Hanley et al., 2023). This deliberate approach contributes to an increase in individuals’ beliefs in misinformation by fostering illusory pattern perception, and promoting associative inference (Van Prooijen et al., 2018; Lee et al., 2023). In the early months of the COVID-19 pandemic, social media platforms were inundated with various conspiracy theories related to the origins, spread, and treatment of the virus (Erokhin et al., 2022). The propagation of misinformation with coordinated conspiracies promotes distrust in public health authorities and skepticism about the safety and effectiveness of COVID-19 vaccines (Sharma et al., 2022; Schmitz et al., 2023). While the proliferation of conspiracy theories on social media has raised significant concerns, little is known about how the incorporation of conspiracy theories in misleading narratives differentiates individuals’ engagement on social media compared to other types of misinformation (Kim and Kim, 2023). This gap calls for exploring how conspiracy theories affect engagement with misinformation during health crises.
Research questions: In this paper, we aim to provide a comprehensive understanding of online engagement with multi-topic misinformation during the early stages of the COVID-19 pandemic, focusing on topic diversity and conspiracy theories. To this end, we investigate the following Research Questions (RQs):
RQ1: Does misinformation exhibit higher topic diversity compared to true information?
RQ2: Does misinformation containing conspiracy theories exhibit higher topic diversity compared to misinformation without conspiracy theories?
RQ3: Does misinformation with higher topic diversity receive more engagement compared to true information?
RQ4: Does misinformation containing conspiracy theories receive more engagement compared to misinformation without conspiracy theories?
Of note, we examine engagement at two levels: news sharer level and post viewer level. News sharers create posts on X that link to source news items outside of the platform. Post viewers engage with news posts within the platform through the reactions of reposting, liking, or replying. The two user groups may have different motivations for engagement (Hanley et al., 2023; Teng et al., 2022; Ballard et al., 2022). Additionally, while previous studies have extensively explored post viewer engagement (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021a; Solovev and Pröllochs, 2022), research on news sharer engagement remains limited. Understanding this under-investigated area is crucial to complement our knowledge of news-sharing behavior.
Methodology: We conduct our analysis based on a COVID-19 dataset consisting of fact-checked source news claims and their corresponding posts on X. These source news claims span six languages and cover the initial period of COVID-19 from January to September 2020. At news sharer level, the engagement metrics include the number of posts linking to a specific news item and the lifetime of the news item, which indicates the duration from the creation of the first post to the latest post. At post viewer level, the engagement metrics include repost count, like count, and reply count. To gain a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we additionally use over 15 million daily sampled posts at the early stage of COVID-19 to train a Word2Vec model and collect topic keywords. Subsequently, we use the topic keywords to calculate topic diversity and identify conspiracy theories. Finally, we employ multiple regression models to investigate our research questions.
Contributions: Our work reveals that elevated topic diversity has been a characteristic of COVID-19 misinformation, particularly for those accompanied by conspiracy narratives. This characteristic becomes more significant within the false news posts on X compared to their source news claims. At post viewer level, the heightened topic diversity and the integration of conspiracy theories significantly correlate with increased engagement (i. e., reposts, likes, and replies) with misinformation on X. For instance, false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. However, at news sharer level, topic diversity and conspiracy theories have no significant association with the post count of source news items. This difference suggests that news sharers and post viewers may have different engagement patterns. Our findings offer valuable insights into understanding the engagement with multi-topic misinformation from news sharers to post viewers on social media during health crises. Additionally, by highlighting the different engagement patterns at news sharer and post viewer levels in terms of topic diversity and conspiracy theories, this research informs the development of targeted interventions and strategies at both user levels to mitigate engagement with false information and foster a more resilient society.
2. Background and Related Work
2.1. Misinformation on Social Media
Social media platforms have transformed information exchange, enabling public users to create and engage with online content globally and instantaneously, albeit without the quality control mechanisms inherent in traditional media. This has led to a significant rise in misinformation. Particularly, the prevalence of misinformation on social media has posed far-reaching consequences across many critical societal domains ranging from politics to public health, becoming a concerning global issue in recent years (Ecker et al., 2022; Pierri et al., 2022; Enders et al., 2022; Green et al., 2022). Consequently, social media platforms are under increasing pressure to implement measures to reduce the spread of misinformation. Understanding how users engage with misinformation and contribute to its viral spread on social media can help social media platforms to design targeted interventions. Prior research has shown that online misinformation spreads farther, faster, deeper, and more broadly compared to true information, as evidenced by the examination of online posts fact-checked by third-party organizations (Vosoughi et al., 2018). Nevertheless, another study, based on community fact-checked posts on X, indicates that misinformation receives fewer reposts compared to true information (Drolsbach and Pröllochs, 2023). This contradiction highlights that not every piece of misinformation spreads faster than true information. Instead, there are likely crucial factors behind the veracity of information that make misinformation unique and contribute to its online diffusion.
Several works have focused on understanding what characterizes misinformation and contributes to its viral spread. For instance, misinformation is more likely to incorporate negative emotions, such as anger, which have been identified as a main driver for information diffusion (Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b). Additionally, there is a partisan asymmetry that misinformation tends to be right-leaning and promote conservative positions (Chuai et al., 2023). Social media users are more likely to believe and share information that is congruent to their political ideologies (Garrett and Bond, 2021; Rao et al., 2022; Robertson et al., 2023a; Osmundsen et al., 2021; Jenke, 2023). Therefore, right-leaning users are more likely to share misinformation on social media, compared to left-leaning users (Nikolov et al., 2021). More importantly, uncertain situations, such as emergency public events or crises, create fertile ground for online misinformation to thrive. During such times, people are eager to know more information about the situations and thus more likely to believe and share related misinformation, which can destroy the trust in the control policies and cause detrimental harm to societies and individuals.
2.2. Misinformation During Public Events/Crises
During public events or crises, characterized by heightened emotions, uncertainty, and a surge in information consumption, falsehoods that cater to specific demands and offer related information to contextualize the situation become rampant and easily spread (Chuai and Zhao, 2022). The preservation of democracy is crucial during political elections. However, democracy can be undermined by the widespread dissemination of dedicated online misinformation aimed at pursuing specific political agendas. It has been found that people are exposed to misinformation more often during elections than usual (Allcott and Gentzkow, 2017). Election-related misinformation is often polarized, which can increase people’s susceptibility and affect their voting behaviors (Rathje et al., 2021; Nikolov et al., 2021; Recuero et al., 2020). Additionally, climate change poses challenges at multiple levels, with any mitigation efforts carrying significant economic and political implications (Lewandowsky, 2021). Consequently, the debate surrounding global climate change has persisted for years and has become increasingly polarized on social media (Fischer et al., 2022; Farrell, 2019; Falkenberg et al., 2022). This polarized environment fosters climate change denial and misinformation, further complicating efforts to address and control climate change (Lewandowsky, 2021; Falkenberg et al., 2022).
The spread of misinformation during health crises is particularly severe, as it can directly harm human health and even cost lives (Chen et al., 2022). Throughout the COVID-19 pandemic, a significant volume of misinformation regarding a wide variety of COVID-19 related topics has circulated widely on social media platforms. For example, the widespread COVID-19 vaccine misinformation reduces users’ beliefs in its effectiveness, impeding the fight against the spread of COVID-19 (Enders et al., 2022; Loomba et al., 2021; Pierri et al., 2022). A global survey has revealed that a substantial part of people consider COVID-19 misinformation as highly reliable (Roozenbeek et al., 2020). Understanding how misinformation spreads and engages users during such public events and crises is crucial for developing effective intervention strategies for emergency management.
2.3. Engagement With COVID-19 Misinformation on Social Media
The engagement with misinformation during health crises, particularly COVID-19, posed significant threats to the pandemic control and public safety. Several studies have scrutinized the factors that can influence engagement with COVID-19 misinformation on social media. Recognizing emotions as a significant driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a), research in this body has predominantly focused on the role of emotions during COVID-19 (Solovev and Pröllochs, 2022; Chuai et al., 2022; Freiling et al., 2021; Charquero-Ballester et al., 2021). For example, Solovev and Pröllochs (2022) conducted a large-scale computational analysis on the role of moral emotions in the spread of COVID-19 misinformation. They found that false posts with higher other-condemning emotions (e. g., anger) are more viral (i. e., more reposts) than true posts. Based on experimental design, Freiling et al. (2021) investigated the role of anxiety in misinformation belief and sharing during COVID-19, revealing that anxiety is a driving factor in believing and sharing both true and false claims. Wang and Rzeszotarski (2023) put the focus on the images and found that online posts containing COVID-19 misinformation images receive similar interactions with posts containing random images. However, COVID-19 misinformation images are shared for longer periods than non-misinformation ones. Additionally, COVID-19 misinformation, mimicking the format and language features of news and scientific reports, is likely to receive more likes (Ngai et al., 2022). For the concern that social bots may share a large amount of misinformation during COVID-19, Teng et al. (2022) analyzed different types of users in the engagement with misinformation. They found that the role of social bots in misinformation sharing is limited, and individuals with human-like behavior play a more prominent role.
Knowing the potential role of topics in engagement with misinformation, Ngai et al. (2022) focused on the specific topics in the engagement with COVID-19 vaccine misinformation. They found that safety concern was the most significant content theme but negatively associated with likes and shares. However, during COVID-19, various aspects of society and individuals became interconnected. COVID-19 misinformation covered a broad spectrum of topics, including public health, politics, medicine, vaccines, and so forth (Wang et al., 2022; Chandrasekaran et al., 2020). The multi-topic misinformation about COVID-19 carries high uncertainty and easily causes confusion to individuals, thus posing threats to people and society (Wang et al., 2022). It has been revealed that individuals can adaptively link multiple related topics and make novel connections through the mechanism of associative inference, which increases the susceptibility to misinformation (Lee et al., 2023).
2.4. Multi-Topic Context and Conspiracy Theories During COVID-19
The discussion on COVID-19 during the pandemic has not only focused on the direct health impacts but has also spilled over into various social, economic, and political domains. Within the social domain, at the beginning of the pandemic, rumors and conspiracy theories about the origin of the virus in China were rampant on social media, which exploits anti-Asian racism and hate speech on social media (Huang et al., 2023). Using COVID-19 as a weapon, malicious misinformation fosters online hate to spread quickly beyond the control of social media platforms (Velásquez et al., 2021). The rising anti-Asian racism during the COVID-19 pandemic knocked Asian businesses (Huang et al., 2023; Luca et al., 2024; Tang et al., 2023). In the economic domain, the pandemic hit the global economy, particularly impacting sectors such as travel and leisure (Wang et al., 2023b). Additionally, the scientific uncertainty surrounding the COVID-19 pandemic fuels the weaponization of COVID-19 for politicization (Kreps and Kriner, 2020). Some individuals and groups have sought to exploit the pandemic to spread extremist ideologies and radicalize others (Davies et al., 2023). The high degree of politicization within COVID-19 information contributes to polarization in attitudes towards the pandemic and even to political ideologies (Hart et al., 2020). For example, the discussion on vaccination become gradually polarized, expanding far beyond the medical discourse and entering the domain of organizational politics (Dehghan and Nagappa, 2022). Taken together, the COVID-19 pandemic is incorporated into many other societal domains to pursue specific interests, yielding significant spillover effects on the stability of societies. In this paper, we use topic diversity to denote the extent to which the information involves multiple topics and explore its impact on the engagement with misinformation on social media during COVID-19.
Particularly, at the onset of the COVID-19 pandemic, a myriad of conspiracy theories emerged, providing alternative explanations for the origins, spread, and treatment of the virus (Erokhin et al., 2022). Examples include claims suggesting that the COVID-19 virus is a man-made bioweapon (Moffitt et al., 2021; Erokhin et al., 2022). Conspiracy theories often revolve around the notion of secret plots orchestrated by malevolent and powerful groups, deliberately crafted to achieve specific outcomes. This characteristic of conspiracy theories facilitates the connection of multiple topics through an illusory perception pattern, potentially fostering associative inference among individuals (Lee et al., 2023; Miani et al., 2022; Van Prooijen et al., 2018). The integration of conspiracy theories within misinformation became more pronounced during the COVID-19 pandemic, marked by a significant increase in hyperlinks from misinformation websites to conspiracy theories (Hanley et al., 2023). Despite this observation, the question of whether misinformation containing conspiracy theories receives more engagement than other types of misinformation remains to be explored. In light of these considerations, this paper tries to address the current research gaps by investigating the roles of topic diversity and conspiracy theories in the engagement with misinformation during health crises.
3. Data and Methods
3.1. Dataset: MM-COVID
The data for our analysis is sourced from a publicly available repository on GitHub (Li et al., 2020). This repository provides an extensive COVID-19 dataset that consists of source news claims and their corresponding engagement on X. Initially, the repository gathers these news claims and their veracity verdicts from reputable third-party fact-checking organizations worldwide, including Snopes and Poynter, as well as official health websites. To keep the quantity of the source claims in each language, the repository filters six languages: English, Spanish, Portuguese, Hindi, French, and Italian, which enables a broad and diverse examination of misinformation across linguistic and cultural contexts. Subsequently, online posts that explicitly reference these news sources are collected from X through its advanced search API. Notably, the news sources in the dataset are linked to the original news URLs. However, due to the diverse range of news items collected from various languages and web pages, precisely parsing the original news pages and obtaining their content is challenging. Instead, the dataset provides news claims, representing the main opinions and topics.
Data preprocessing: We begin by translating all the non-English texts in source news claims and their associated posts into English using the Googletrans Python package. Subsequently, we perform text preprocessing by removing non-alphabetic and non-numerical characters, stop words, and URLs. We lemmatize all the words using the English transformer pipeline on Spacy and only consider news claims and posts that contain at least one lemmatized word. Additionally, we exclude posts that were posted before January 1, 2020. Due to the absence of release dates for source news items in the dataset, we use the first post that links to the specific source news item as its release date. In the end, we obtain news claims and associated posts. Specifically, posts are linked to false news claims, and posts are linked to true news claims.
3.2. COVID-19 Topic Spectrum
A previous study analyzed over 13 million English COVID-19 posts and identified 26 topics within 10 broad themes, including source, prevention, spread and growth, treatment and recovery, impact on the economy and markets, impact on health care sector, and government response, political impact, and racism (Chandrasekaran et al., 2020). This research also provided keywords corresponding to each topic. To ensure a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we utilize these keywords as seeds to identify similar words and construct a topic lexicon based on a Word2Vec model. Notably, after reexamination, we remove the topic about locations that have no specific meanings and consider the remaining 25 topics across 10 themes (Table 1).
Theme | Topic |
Source (origin) | Outbreak, alternative causes |
Prevention | Social distancing, disinfecting and cleanliness |
Symptoms | Symptoms |
Spread and growth | Modes of transmission, spread of cases, death reports |
Treatment and recovery | Drugs and vaccines, therapies, alternative methods, |
testing | |
Impact on health care sector | Impact on hospitals and clinics, policy changes, |
frontline workers | |
Government response | Travel restrictions, financial measures, |
lockdown regulations | |
Impact on the economy and markets | Shortage of products, panic buying, stock markets, |
employment, impact on business | |
Political impact | Political impact |
Racism | Racism |
Word2Vec model: To train the Word2Vec model, we additionally collect a large-scale COVID-19 dataset from a public repository on Kaggle.111https://www.kaggle.com/datasets/smid80/coronavirus-covid19-tweets This dataset contains posts in multiple languages, posted from March 9 to April 30, 2020 on X. Given the size of the dataset, we only filter English posts and do not involve the translation process. By combining this dataset with MM-COVID, we obtain a total of English posts. After removing non-alphabetic and non-numerical characters as well as URLs in the posts, we utilize the refined corpus to train a Word2vec model with a vector size of 200 using the gensim library.222The model is trained on the High Performance Computing platform at the University of Luxembourg. The examples of the keywords and their ten most similar words are shown in Table 2.
bioweapon | vaccine | election | quarantine |
biowarfare | vaccines | elections | isolation |
biologicalweapon | vaccination | primaries | quarantined |
bioweapons | vaccin | voting | selfquarantine |
bioengineered | vax | 2020election | quarentine |
biologicalwarfare | vacine | electoral | selfisolation |
bioterrorism | cure | election2020 | quarantining |
engineered | inoculation | referendum | qurantine |
weaponized | treatments | vote | lockdown |
wuhanlab | coronavirusvaccine | votes | quarintine |
biowar | antibodies | generalelection | isolating |
Topic lexicon: We expand our set of topic keywords by incorporating words that exhibit similarities of more than 0.6 with the initial topic seeds. This process yields 828 additional topic keywords. Next, we conduct a manual review of these keywords, identifying that 90.6% of them are indeed relevant to the corresponding topic seeds and effectively represent the associated topics. Consequently, we integrate the topic seeds with these related keywords to create a comprehensive topic lexicon.
Topic identification: We utilize the topic lexicon to identify topics discussed in both news claims and their associated posts. As a result, 82.4% of news claims contained at least one topic. Additionally, our analysis considers both the content of the source claims and the content of the corresponding posts when identifying topics in online posts. Consequently, 91.4% of posts contain at least one topic.
3.3. Topic Diversity
Topic diversity indicates the extent to which the content involves multiple topics. To give a sense of topic diversity, we construct topic co-occurrence networks for true and false news claims. Specifically, we calculate the frequency of topic co-occurrence of any two topics out of 25 topics in true and false news claims, and retain the co-occurrence edges with frequencies exceeding the median threshold. In the network of true news claims (Fig. 1(a)), the number of edges is 107, while in the network of false news claims (Fig. 1(b)), the number of edges is 167. Compared to true news claims, false news claims contain 56.1% more co-occurrence edges and are more interconnected, indicating higher interconnectedness and suggesting greater topic diversity in false news claims. Building upon previous research that used the Gini coefficient to measure the inequality of topic distribution in the content (Miani et al., 2022), we quantitatively measure topic diversity () as:
where denotes the number of words associated to topic in each news claim (post), and is the number of topics. In our analysis, we exclude news items and posts that do not relate to any topic, resulting in source news claims with corresponding posts (see details in Table 4).
3.4. COVID-19 Conspiracy Theories
Conspiracy theories have been often thought of as a subset of misinformation (Hanley et al., 2023). Following the method commonly adopted by previous related research (Erokhin et al., 2022; Moffitt et al., 2021; Sharma et al., 2022), we select a set of keywords (phrases) from the topic lexicon to identify the conspiracy theories embedded in the false news items and associated posts. Specifically, we consider eight COVID-19 conspiracy theory categories that connect to at least one topic in our analysis: the sources and origins of the virus that caused the outbreak of COVID-19 (VirusOrigin), the connection between 5G and COVID-19 (5G), the role of big pharmaceutical companies (BigPharma) and Bill Gates (BillGates) in the spread of COVID-19, the scope of the pandemic is exaggerated (Exaggeration), the claim that hospitals were actually empty (FilmYourHospital), the idea that the genetically modified crops led to the emergence of COVID-19 (GMO), and skepticism regarding the ineffectiveness and side-effects of vaccines (Vaccines).
As shown in Table 3, we match the keywords with the lemmatized words in each news claim or post, and for phrases, we directly match them with the original translated text. We find that false news claims and false news posts incorporate keywords (phrases) indicating conspiracy theories. Given that the share of conspiracy posts ranges from 0.6% to 18% (Erokhin et al., 2022), the 6% share of posts containing conspiracy theories in our dataset is reasonable. Upon a manual check of news items that contain keywords or phrases related to conspiracy theories, we find that the accuracy is 92.1%, which suggests that the selected keywords (phrases) can effectively identify the conspiracy theories integrated into the misinformation (Erokhin et al., 2022; Moffitt et al., 2021).
Conspiracy theory | Keywords |
Origin | bat, batsoup, weapon, biowar, lab, conspiracy, chinese()virus, |
biochemical, bioterrorism, biowarfare, china()virus, wildlife, ccp()virus, | |
chinese()flu, china()coronavirus, chinese()coronavirus, chinesevirus19, | |
chines()virus, manmade, wuhan()coronavirus, wuhan()coronavius, | |
wuflu,wuhan()virus, chinaliedpeopledie | |
Vaccines | {vaccine, vax, vaccination, vaccin, vacine} and {anti, infertile, |
do not work, don’t work, does not work, doesn’t work, autism, | |
autoimmune} | |
5G | 5G |
BigPharma | big pharma, fauci pharma, gates pharma |
BillGates | Bill Gates |
Exaggeration | does not exist, doesn’t exist, exaggerated, inflated, refuse |
FilmYourHospital | filmyourhospital, film your hospital, empty hospital, empty bed |
GMO | gmo, genetically modified |
3.5. The Engagement From Source News to Posts on X
On the platform of X, users can interact with posts they view through reposting, liking, or replying. Existing research often uses the number of reposts to study engagement with misinformation on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b; Drolsbach and Pröllochs, 2023; Chuai et al., 2024). Additionally, the number of likes and the number of replies are also important indicators of engagement with online posts (Chuai et al., 2024; Papakyriakopoulos and Goodman, 2022). Notably, news items outside of X enter the social environment through news sharers on X. Multiple posts on X can link to the same original news items, which increases the visibility of the news items. However, few studies examine engagement with misinformation from the perspective of source news items to posts. In this paper, we consider user engagement at two levels: news sharer level and post viewer level (Fig. 2). At news sharer level, news items are shared via posts on X. The engagement metrics include the number of posts and the lifetime of news items from the creation of the first post to the creation of the latest post. At post viewer level, engagement is measured by how viewers interact with these posts. Post viewers can repost, like, and reply to the posts that are created by news viewers and contain links to the original news items.
3.6. Empirical Models
We employ multiple explanatory regression models to explore the research questions separately. To this end, we define the following dependent variables to measure engagement:
: The duration (in hours) from the creation of the first post to the creation of the latest post on X for the specific news item.
: A count variable indicating the number of posts associated with the specific news item.
: A count variable indicating the number of reposts received by the specific post.
: A count variable indicating the number of likes received by the specific post.
: A count variable indicating the number of replies received by the specific post.
The key independent variables include:
: A dummy variable indicating whether the news claim is false ( 1) or true ( 0).
: A dummy variable indicating whether the false news claim or post contains conspiracy theories ( 1) or not ( 0).
: The topic diversity of the news claim or post.
We also control for the variables from post and user characteristics:
: A continuous-like count variable indicating the number of words in the text of the news claim or post.
: A dummy variable indicating whether the post has media elements ( 1) or not ( 0).
: A dummy variable indicating whether the post account is verified ( 1) or not ( 0).
: A continuous-like count variable indicating the age (in days) of the post account.
: A continuous-like count variable indicating the number of followers of the post account.
: A continuous-like count variable indicating the number of followees of the post account.
Additionally, emotions are considered as a main driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a). Our analysis considers six cross-cultural emotions – anger, disgust, fear, joy, sadness, and surprise – out of the eight basic emotions (Chuai and Zhao, 2022; Plutchik, 1991). We employ the commonly used NRC emotion lexicon to count emotion words in news claims and posts. Notably, 92.1% news claims and 95% posts contain at least one emotion word. Subsequently, we calculate the weight of each emotion in the specific text content based on the number of corresponding emotion words. If the news claims (posts) have no emotion words, all the weights of emotions are set to 0. The descriptive statistics for the variables are shown in Table 4.
(1) | (2) | |
All | CTs | |
News Count | 7,273 | 82 |
Falsehood | 18.9% | 100% |
Engagement metrics | ||
Lifetime | 315.617 (900.617) | 1,968.141 (1,626.337) |
Post Count | 70,904 | 4,265 |
Repost Count | 1,300,597 | 161,823 |
Like Count | 3,943,848 | 452,106 |
Reply Count | 380,224 | 107,363 |
News characteristics | ||
Dates | 01/02/20 – 09/14/20 | 01/07/20 – 09/08/20 |
Diversity | 0.083 (0.045) | 0.089 (0.038) |
Words | 15.614 (6.364) | 11.841 (6.739) |
Anger | 0.058 (0.109) | 0.040 (0.093) |
Disgust | 0.045 (0.108) | 0.059 (0.153) |
Fear | 0.183 (0.208) | 0.149 (0.183) |
Joy | 0.065 (0.128) | 0.097 (0.215) |
Sadness | 0.096 (0.139) | 0.089 (0.135) |
Surprise | 0.045 (0.099) | 0.026 (0.080) |
Post characteristics | ||
Dates | 01/02/20 – 09/14/20 | 01/09/20 – 09/09/20 |
Diversity | 0.104 (0.052) | 0.111 (0.050) |
Words | 26.372 (10.192) | 27.006 (8.876) |
Anger | 0.068 (0.100) | 0.069 (0.090) |
Disgust | 0.074 (0.134) | 0.064 (0.105) |
Fear | 0.195 (0.177) | 0.157 (0.146) |
Joy | 0.062 (0.106) | 0.088 (0.163) |
Sadness | 0.129 (0.133) | 0.095 (0.106) |
Surprise | 0.052 (0.091) | 0.050 (0.088) |
Media | 21.9% | 13.6% |
Verified | 18.5% | 7.4% |
AccountAge | 2,460.659 (1,457.714) | 2,259.347 (1,393.434) |
Followers | 369,041.439 (1,755,545.224) | 56,847.547 (1,399,857.795) |
Followees | 3,031.517 (15,099.135) | 2,361.965 (15,535.361) |
Topic diversity (RQ1 & RQ2): We use as the dependent variable to analyze the topic diversity in true and false news from original news claims to online posts. In terms of the claims, the linear regression model that incorporates the main independent variables of and is specified as:
where indicates six basic emotions. Additionally, denotes language-specific fixed effects, and denotes month-year fixed effects.
Given that each news item can have multiple posts, we need to control for the possible heterogeneity among news items when analyzing the topic diversity in terms of online posts. Therefore, we specify a liner regression model incorporating news-specific random effects:
where indicates news-specific random effects. Additionally, we also incorporate , poster characteristics as control variables in the model.
Engagement at news sharer level – lifetime and post count (RQ3 & RQ4): The first layer of engagement is from source news claims to news posts on X. We examine how topic diversity and conspiracy theories influence engagement with misinformation in terms of news lifetime and post count. First, we take as the dependent variable and specify a linear regression model as follows:
where the interaction term of is to estimate the moderating effect of on . Subsequently, we use a negative binomial regression model to explain . The dependent variable is , and the independent variables remain consistent with Eq. (4).
Engagement at post viewer level – reposts, likes, and replies (RQ3 & RQ4): To examine the effects of topic diversity and conspiracy theories on social engagement (i. e., reposts, likes, and replies) with misinformation on X/Twitter, we specify a negative binomial regression model with language-specific fixed effects, month-year fixed effects, and news-specific random effects:
where indicates all the independent variables, and signifies the dependent variables, including , and . All the continuous variables in the above models are z-standardized.
4. Empirical Results
4.1. Topic Diversity (RQ1 & RQ2)
Summary statistics: Fig. 3(a) shows that false news claims (mean of 0.086) exhibit slightly higher topic diversity compared to true news claims (mean of 0.082, , ). Notably, the difference in topic diversity between true and false news items is magnified at the post level, where the mean topic diversity in false news posts is 0.113, compared to 0.079 in true news posts (Fig. 3(b), , ). Additionally, from claims to posts (Fig. 3(c)), topic diversity in false news items (mean of 0.031) increases more than that in true news items (mean of 0.012, , ). For false news items containing conspiracy theories (CTs), the topic diversity at the claim level has no statistically significant difference compared to the overall false news items (Fig. 3(a), , ). However, Fig. 3(c) indicates a higher proportion of the segment with most significant increases in topic diversity for CT news items, compared to false news items ( , ), which may lead to the significant difference in topic diversity between CT news and general false news at the post level (Fig. 3(b), , ).
Regression results: In Column (1) of Table 5, the coefficient estimate of is significantly positive ( , ). This implies that false news exhibits higher topic diversity compared to true news at the claim level, aligning with the observations in Fig. 3(a). Additionally, the coefficient estimate of is not statistically significant, which indicates that the topic diversity in false news claims containing conspiracy theories is not significantly different from that in false news claims without conspiracy theories. Subsequently, we explore the topic diversity between true and false news at the post level. In Column (2) of Table 5, the coefficient estimate of is significantly positive ( , ), suggesting that false news maintains higher topic diversity than true news at the post level.
(1) | (2) | (3) | |
News | Post | News Post | |
Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | |
Verified | ∗∗∗ | ||
AccountAge | |||
Followers | ∗∗∗ | ||
Followees | ∗∗∗ | ∗∗∗ | |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✗ | ✓ | ✓ |
0.179 | ✗ | ✗ |
Notably, as observed in Fig. 3(b) and the coefficient estimates of in Columns (1) and (2), the difference between true and false news posts is larger than that between true and false news claims. Additionally, the coefficient estimate of becomes significantly positive ( , ), which suggests that false news posts containing conspiracy theories exhibit even higher topic diversity than false news posts without conspiracy theories. There are two possible reasons for the larger difference in topic diversity between true and false news at the post level compared to the claim level. One is that news claims with higher topic diversity receive more posts on X. Another is that false news has a bigger increase in topic diversity from the claim level to the post level, compared to true news. Given the second reason, we further examine the change in topic diversity from the source news claims to the corresponding posts. In Column (3) of Table 5, the positive coefficient estimate of is statistically significant ( , ). It suggests that the increase of topic diversity in false news from the claim level to the post level is bigger than that in true news, which could partially explain the enlarged difference in topic diversity between true and false news posts, compared to their source claims. In addition, the coefficient estimate of in Column (3) of Table 5 is also significantly positive ( , ). It means that the topic diversity in false news that contains conspiracy theories increases more than that in false news without conspiracy theories from the claim level to the post level. This can explain why the coefficient estimate of in Column (1) of Table 5 is statistically insignificant, whereas it becomes statistically significant in Column (2).
4.2. News Sharer Engagement: Lifetime and Post Count (RQ3 & RQ4)
Summary statistics: Fig. 3(d) shows that false news items have significantly longer lifetimes (mean of 1463.757 hours) than true news items (mean of 47.951 hours, , ). CT news items have longer lifetimes (mean of 1968.141 hours) than false news items ( , ). Additionally, Fig. 3(e) shows that false news claims receive more posts (mean of 38.015) than true news claims (mean of 3.159, , ). However, the number of posts received by CT news claims (mean of 44.976) has no statistically significant difference from that received by false news claims ( , ).
Regression results: As shown in Column (1) of Table 6, the coefficient estimate of is significantly positive ( , ), which suggests that false news items have longer lifetimes to get new posts on X than true news items. Additionally, the coefficient estimate of is also significantly positive ( , ). It means that false news items containing conspiracy theories have significantly longer lifetimes than false news items without conspiracy theories. However, the coefficient estimates for and are statistically insignificant. This suggests that topic diversity has no significant impact on the lifetime of news items, irrespective of the veracity.
Additionally, in Column (2) of Table 6, the coefficient estimate of is significantly positive ( , ), indicating that the number of posts referencing a specific source news claim is significantly associated with its veracity. On average, false news claims are linked by times more posts on X compared to true news claims. Notably, the coefficient estimate of is not statistically significant for post count, which implies that the integration of conspiracy theories has no additional impact on the number of posts received by false news claims. The coefficient estimate of is statistically insignificant. This suggests that the topic diversity of a source news claim has no significant relationship with the number of online posts it receives. The reason for the enlarged difference in topic diversity between true and false news posts is not supported by this observation. Furthermore, the coefficient estimate for the interaction is also statistically insignificant, reinforcing the notion that topic diversity has no significant effect on post count, irrespective of the veracity of news claims.
(1) | (2) | |
Lifetime | Post Count | |
Falsehood | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗ | |
Diversity | ||
Falsehood Diversity | ||
Words | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ |
Language | ✓ | ✓ |
MonthYear | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ |
0.481 | 0.188 |
4.3. Post Viewer Engagement: Reposts, Likes, and Replies (RQ3 & RQ4)
Summary statistics: Figs. 3(f) – 3(h) reveal that, on average, true news posts get more engagement on X than false news posts in terms of reposts (true: 29.396, false: 14.403, , ), likes (true: 84.765, false: 45.234, , ), and replies (true: 6.294, false: 5.031, , ). However, it’s worth noting that the news posts that receive the most significant reposts, likes, or replies belong to the false category. Additionally, CT news posts receive more reposts (mean of 37.942, , ), likes (mean of 106.004, , ), and replies (mean of 25.173, , ) than overall false news posts. Notably, the difference in reply count between true and false news posts is the smallest compared to repost count and like count.
Regression results: In Column (1) of Table 7, the coefficient estimate of is significantly negative ( , ), which means that false news posts receive fewer reposts than true news posts. On average, false news posts receive 22.6% fewer reposts than true news posts. Similarly, the coefficient estimate of in Column (2) is significantly negative ( , ). On average, false news posts receive 14.4% fewer likes than true news posts. However, the coefficient estimate of in Column (3) is significantly positive ( , ). This means that, on average, false news posts receive 49.2% more replies than true news posts. Together with the observations in Figs. 3(f) – 3(h), the results indicate that the interaction of reply with the veracity of posts may vary from the interactions of reposting and liking.
Additionally, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive. It indicates that heightened topic diversity is associated with increased online engagement in terms of reposts, likes, and replies. However, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly negative. This suggests that the positive association of topic diversity with engagement on X is reduced within the context of false news posts. On the contrary, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive. This means that the integration of conspiracy theories is positively linked to the engagement with false news posts on X at post viewer level. Specifically, false news posts that contain conspiracy theories receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
(1) | (2) | (3) | |
Repost Count | Like Count | Reply Count | |
Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✓ | ✓ | ✓ |
4.4. Robustness Checks
To ensure the robustness of our findings, we conduct several additional checks. (i) The cross-correlations among the independent variables are reported in Table S1 (news level) and Table S2 (post level). The variance inflation factors (VIFs) are also reported in Table S3. All the VIFs are significantly lower than the threshold of 4, which indicates that multicollinearity is not an issue for our analysis. (ii) We recalculate topic diversity at the theme level and repeat our analysis (Table S4 – Table S6). (iii) We further balance the samples between true and false news items through propensity score matching and reduce potential unobserved confounding factors (Table S7 – Table S9). (iiii) We use an alternative topic modeling method (BERTopic) to cluster topics again and repeat our analysis (Table S10 – Table S12). All the results are robust and consistently support our main findings.
5. Discussion
The spread of misinformation surrounding health crises is a serious public concern. In this paper, we present a comprehensive analysis of the impact of topic diversity and conspiracy theories on the engagement dynamics of COVID-19 misinformation. Specifically, we examine user engagement at two levels spanning from the origin as source news items to their propagation through online posts and subsequent social engagement (i. e., reposts, likes, and replies) on X. Our analysis yields several key findings: (i) False news, especially when accompanied by conspiracy theories, exhibits higher topic diversity compared to true news. This discrepancy increases from source news claims to news posts on X. (ii) At news sharer level, neither the news lifetime nor the number of news posts is significantly associated with topic diversity. However, false news has a longer lifetime and receives more posts compared to true news. Additionally, the integration of conspiracy theories is associated with a longer lifetime for false news but has no significant association with the number of false news posts. (iii) At post viewer level, posts with higher topic diversity receive more reposts, likes, and replies. However, the positive effect of topic diversity is moderated by the news veracity, with higher topic diversity being linked to more engagement with true news posts compared to false news posts. (iiii) The integration of conspiracy theories is linked to more engagement with COVID-19 misinformation at post viewer level. False news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. These findings have useful implications for both research and practical applications.
Implications: First, the observed high topic diversity in COVID-19 misinformation indicates that the interconnection of multiple topics becomes a characteristic of misinformation narratives during health crises. Previous research also reveals a coordinated effort among misinformation websites from multiple domains to hyperlink to each other (Hanley et al., 2023). Notably, the topic diversity in COVID-19 misinformation significantly increases from its source claims to online posts. This mutation may contribute to the adaptability of misinformation, allowing it to resonate with diverse audiences and contexts (Yan et al., 2022). Additionally, the finding that misinformation containing conspiracy theories exhibits the strongest topic diversity aligns with previous research (Miani et al., 2022), highlighting the interconnected nature of conspiracy narratives. Within the context of multi-topic misinformation, individuals tend to adaptively connect related information and make associative inferences to reduce uncertainty and gain control in a way that is congruent to their prior knowledge (Lee et al., 2023; Miani et al., 2022). Understanding this cognitive process is essential for developing targeted interventions aimed at reducing susceptibility to multi-topic misinformation, especially when accompanied by conspiracy theories.
Second, at the news sharer level, the observed longer lifetime and higher post count on X for false news items, compared to true news items, indicate the endurance and prevalence of misinformation on social media during COVID-19. Notably, few prior studies focus on engagement at news sharer level and study how false news enters the social media platforms via news sharers. The extended lifetime we observe in this paper may provide false news with more opportunities to accumulate engagement on social media, potentially reaching a broader audience. Moreover, the lifetime and post count appear to be largely attributed to the veracity of the news claims, as opposed to other variables including and . Given this, it is plausible that the prolonged lifetime and high post count for false news claims mainly result from deliberately planned dissemination strategies by misinformation providers, involving means such as social bots and monetization (Teng et al., 2022; Moffitt et al., 2021; Ballard et al., 2022). Consequently, the role of individuals’ self-motivated sharing behaviors in the endurance and prevalence of COVID-19 misinformation is limited at news sharer level. However, massive posts repeating the same false claim can still increase individuals’ susceptibility and beliefs. In this regard, social media platforms should take measures to restrict repeated original posts from the same false claims and their mutations, which is a lack of discussion in the previous research.
Third, as opposed to the engagement at news sharer level, the engagement at post viewer level (i. e., reposts, likes, and replies) on X might be largely driven by human users rather than social bots during COVID-19 (Teng et al., 2022). Organic user engagement from human users can be influenced by multiple factors. For instance, our results in Table 7 show that the engagement at the post viewer level can be affected by factors related to content (e. g., topic diversity, conspiracy theories, emotions) and context (e. g., verified status and follower count). Previous findings on the virality of misinformation at post viewer level are mixed (Vosoughi et al., 2018; Drolsbach and Pröllochs, 2023). We also find that false news posts receive fewer reposts and likes but receive more replies, compared to true news posts. Of note, the veracity of information is not explicit to users before being verified. Therefore, there might be other content or context factors that are related to the veracity and foster engagement. While a large body of research has examined the engagement at the posts viewer level (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b, a; Wang et al., 2023a; Solovev and Pröllochs, 2022), the understanding on the role of topic diversity and conspiracy theories in shaping engagement of post viewers with misinformation is still missing. Our findings show that higher topic diversity is associated with more reposts, likes, and reposts for both true and false COVID-19 news posts on X. However, despite COVID-19 misinformation having higher topic diversity, the positive effect of topic diversity on its engagement is reduced, requiring further detailed study. Notably, conspiracy theories can significantly enhance engagement with COVID-19 misinformation. To efficiently reduce social engagement with (health) misinformation, social media platforms should exclusively focus on the typical misinformation that exhibits high topic diversity and contains conspiracy theories.
Limitations: Our empirical study reports associations between variables and avoids making causal claims. Future research could benefit from experimental designs to establish causal relationships and better understand the influence of topic diversity and conspiracy theories on engagement with online misinformation during health crises. Moreover, our analysis relies on data from X, limiting the generalizability of our findings to other social media platforms. Though the prevalence of health misinformation was the highest on X during COVID-19 (Suarez-Lledo and Alvarez-Galvez, 2021), different platforms may exhibit unique engagement patterns and user behaviors, necessitating further research across the platforms to capture a comprehensive understanding of online misinformation dynamics during health crises. Additionally, the absence of detailed engagement timelines at post viewer level in our dataset hinders a more granular analysis of how topic diversity and conspiracy theories influence engagement with COVID-19 misinformation over time on social media.
6. Conclusion
The spread of health misinformation on social media, especially during health crises such as COVID-19, poses unprecedented threats to our society. It is crucial to understand the factors and mechanisms that contribute to the engagement with health misinformation on social media. Here, we provide a comprehensive analysis of how topic diversity and conspiracy theories shape engagement with COVID-19 misinformation from news sharer level to post viewer level. We find that high topic diversity and the integration of conspiracy theories are characteristics of COVID-19 misinformation. Topic diversity has no significant association with the engagement at news sharer level. However, topic diversity and conspiracy theories can significantly enhance engagement with COVID-19 misinformation at post viewer level. Our findings provide insights into understanding the engagement with health misinformation and highlight the engagement patterns for news sharers and post viewers regarding topic diversity and conspiracy theories. These insights are valuable for developing targeted interventions at both user levels during health crises.
7. Ethics Statement
This research has received ethical approval from the Ethics Review Panel of the University of Luxembourg (ref. ERP 23-053 REMEDIS). All analyses are based on publicly available data. We declare no competing interests.
Acknowledgements.
This research is supported by the Luxembourg National Research Fund (FNR) and Belgian National Fund for Scientific Research (FNRS), as part of the project REgulatory Solutions to MitigatE DISinformation (REMEDIS), grant ref. INTER_FNRS_21_16554939_REMEDIS.References
- (1)
- Allcott and Gentzkow (2017) Hunt Allcott and Matthew Gentzkow. 2017. Social media and fake news in the 2016 election. Journal of Economic Perspectives 31, 2 (2017), 211–236.
- Ballard et al. (2022) Cameron Ballard, Ian Goldstein, Pulak Mehta, Genesis Smothers, Kejsi Take, Victoria Zhong, Rachel Greenstadt, Tobias Lauinger, and Damon McCoy. 2022. Conspiracy brokers: Understanding the monetization of YouTube conspiracy theories. In WWW.
- Chandrasekaran et al. (2020) Ranganathan Chandrasekaran, Vikalp Mehta, Tejali Valkunde, and Evangelos Moustakas. 2020. Topics, trends, and sentiments of tweets about the COVID-19 pandemic: Temporal infoveillance study. Journal of Medical Internet Research 22, 10 (2020), e22624.
- Charquero-Ballester et al. (2021) Marina Charquero-Ballester, Jessica G Walter, Ida A Nissen, and Anja Bechmann. 2021. Different types of COVID-19 misinformation have different emotional valence on Twitter. Big Data & Society 8, 2 (2021), 20539517211041279.
- Chen et al. (2022) Canyu Chen, Haoran Wang, Matthew Shapiro, Yunyu Xiao, Fei Wang, and Kai Shu. 2022. Combating health misinformation in social media: Characterization, detection, intervention, and open issues. ArXiv.
- Chen et al. (2023) Lucia L. Chen, Steven R. Wilson, Sophie Lohmann, and Daniela V. Negraia. 2023. What are you anxious about? Examining subjects of anxiety during the COVID-19 pandemic. In ICWSM.
- Chuai et al. (2022) Yuwei Chuai, Yutian Chang, and Jichang Zhao. 2022. What really drives the spread of COVID-19 tweets: A revisit from perspective of content. In DSAA.
- Chuai et al. (2024) Yuwei Chuai, Haoye Tian, Nicolas Pröllochs, and Gabriele Lenzini. 2024. Did the roll-out of Community Notes reduce engagement with misinformation on X/Twitter?. In CSCW.
- Chuai and Zhao (2022) Yuwei Chuai and Jichang Zhao. 2022. Anger can make fake news viral online. Frontiers in Physics 10 (2022), 970174.
- Chuai et al. (2023) Yuwei Chuai, Jichang Zhao, Nicolas Pröllochs, and Gabriele Lenzini. 2023. Political Elites in False Statements on the Internet. ArXiv.
- Davies et al. (2023) Garth Davies, Edith Wu, and Richard Frank. 2023. A witch’s brew of grievances: The potential effects of COVID-19 on radicalization to violent extremism. Studies in Conflict & Terrorism 46, 11 (2023), 2327–2350.
- Dehghan and Nagappa (2022) Ehsan Dehghan and Ashwin Nagappa. 2022. Politicization and radicalization of discourses in the alt-tech ecosystem: A case study on Gab Social. Social Media + Society 8, 3 (2022), 20563051221113075.
- Douglas and Sutton (2018) Karen M. Douglas and Robbie M. Sutton. 2018. Why conspiracy theories matter: A social psychological analysis. European Review of Social Psychology 29, 1 (2018), 256–298.
- Drolsbach and Pröllochs (2023) Chiara Drolsbach and Nicolas Pröllochs. 2023. Diffusion of community fact-checked misinformation on Twitter. In CSCW.
- Ecker et al. (2022) Ullrich KH Ecker, Stephan Lewandowsky, John Cook, Philipp Schmid, Lisa K Fazio, Nadia Brashier, Panayiota Kendeou, Emily K Vraga, and Michelle A Amazeen. 2022. The psychological drivers of misinformation belief and its resistance to correction. Nature Reviews Psychology 1, 1 (2022), 13–29.
- Enders et al. (2022) Adam M. Enders, Joseph Uscinski, Casey Klofstad, and Justin Stoler. 2022. On the relationship between conspiracy theory beliefs, misinformation, and vaccine hesitancy. Plos One 17, 10 (2022), e0276082.
- Erokhin et al. (2022) Dmitry Erokhin, Abraham Yosipof, and Nadejda Komendantova. 2022. COVID-19 conspiracy theories discussion on Twitter. Social Media + Society 8, 4 (2022), 20563051221126051.
- Falkenberg et al. (2022) Max Falkenberg, Alessandro Galeazzi, Maddalena Torricelli, Niccolò Di Marco, Francesca Larosa, Madalina Sas, Amin Mekacher, Warren Pearce, Fabiana Zollo, Walter Quattrociocchi, and Andrea Baronchelli. 2022. Growing polarization around climate change on social media. Nature Climate Change (2022), 1–8.
- Farrell (2019) Justin Farrell. 2019. The growth of climate change misinformation in US philanthropy: evidence from natural language processing. Environmental Research Letters 14, 3 (2019), 034013.
- Fischer et al. (2022) Helen Fischer, Markus Huff, and Nadia Said. 2022. Polarized climate change beliefs: No evidence for science literacy driving motivated reasoning in a U.S. national study. American Psychologist 77 (2022), 822–835.
- Freiling et al. (2021) Isabelle Freiling, Nicole M Krause, Dietram A Scheufele, and Dominique Brossard. 2021. Believing and sharing misinformation, fact-checks, and accurate information on social media: The role of anxiety during COVID-19. New Media & Society (2021), 14614448211011451.
- Garrett and Bond (2021) R. Kelly Garrett and Robert M. Bond. 2021. Conservatives’ susceptibility to political misperceptions. Science Advances 7, 23 (2021), eabf1234.
- Green et al. (2022) Jon Green, William Hobbs, Stefan McCabe, and David Lazer. 2022. Online engagement with 2020 election misinformation and turnout in the 2021 Georgia runoff election. Proceedings of the National Academy of Sciences 119, 34 (2022), e2115900119.
- Hanley et al. (2023) Hans W. A. Hanley, Deepak Kumar, and Zakir Durumeric. 2023. A golden age: Conspiracy theories’ relationship with misinformation outlets, news media, and the wider Internet. In CSCW.
- Hart et al. (2020) P. Sol Hart, Sedona Chinn, and Stuart Soroka. 2020. Politicization and polarization in COVID-19 news coverage. Science Communication 42, 5 (2020), 679–697.
- Horner et al. (2021) Christy Galletta Horner, Dennis Galletta, Jennifer Crawford, and Abhijeet Shirsat. 2021. Emotions: The unexplored fuel of fake news on social media. Journal of Management Information Systems 38, 4 (2021), 1039–1066.
- Huang et al. (2023) Justin T. Huang, Masha Krupenkin, David Rothschild, and Julia Lee Cunningham. 2023. The cost of anti-Asian racism during the COVID-19 pandemic. Nature Human Behaviour 7, 5 (2023), 682–695.
- Jenke (2023) Libby Jenke. 2023. Affective polarization and misinformation belief. Political Behavior (2023), 1–60.
- Kim and Kim (2023) Soojong Kim and Jisu Kim. 2023. The information ecosystem of conspiracy theory: Examining the QAnon narrative on Facebook. (2023).
- Kou et al. (2017) Yubo Kou, Xinning Gui, Yunan Chen, and Kathleen Pine. 2017. Conspiracy talk on social media: Collective sensemaking during a public health crisis. In CSCW.
- Kreps and Kriner (2020) S. E. Kreps and D. L. Kriner. 2020. Model uncertainty, political contestation, and public trust in science: Evidence from the COVID-19 pandemic. Science Advances 6, 43 (2020), eabd4563.
- Lee et al. (2023) Sian Lee, Haeseung Seo, Dongwon Lee, and Aiping Xiong. 2023. Associative inference can increase people’s susceptibility to misinformation. In ICWSM.
- Lewandowsky (2021) Stephan Lewandowsky. 2021. Climate change disinformation and how to combat it. Annual Review of Public Health 42 (2021), 1–21.
- Li et al. (2020) Yichuan Li, Bohan Jiang, Kai Shu, and Huan Liu. 2020. MM-COVID: A multilingual and multimodal data repository for combating COVID-19 disinformation. ArXiv.
- Loomba et al. (2021) Sahil Loomba, Alexandre de Figueiredo, Simon J. Piatek, Kristen de Graaf, and Heidi J. Larson. 2021. Measuring the impact of COVID-19 vaccine misinformation on vaccination intent in the UK and USA. Nature Human Behaviour 5, 3 (2021), 337–348.
- Luca et al. (2024) Michael Luca, Elizaveta Pronkina, and Michelangelo Rossi. 2024. The evolution of discrimination in online markets: How the rise in anti-Asian bias affected Airbnb during the pandemic. Marketing Science (2024).
- Miani et al. (2022) Alessandro Miani, Thomas Hills, and Adrian Bangerter. 2022. Interconnectedness and (in)coherence as a signature of conspiracy worldviews. Science Advances 8, 43 (2022), eabq3668.
- Moffitt et al. (2021) J. D. Moffitt, Catherine King, and Kathleen M. Carley. 2021. Hunting conspiracy theories during the COVID-19 pandemic. Social Media + Society 7, 3 (2021), 20563051211043212.
- Ngai et al. (2022) Cindy Sing Bik Ngai, Rita Gill Singh, and Le Yao. 2022. Impact of COVID-19 vaccine misinformation on social media virality: Content analysis of message themes and writing strategies. Journal of Medical Internet Research 24, 7 (2022), e37806.
- Nikolov et al. (2021) Dimitar Nikolov, Alessandro Flammini, and Filippo Menczer. 2021. Right and left, partisanship predicts (asymmetric) vulnerability to misinformation. Harvard Kennedy School Misinformation Review (2021).
- Osmundsen et al. (2021) Mathias Osmundsen, Alexander Bor, Peter Bjerregaard Vahlstrup, Anja Bechmann, and Michael Bang Petersen. 2021. Partisan polarization is the primary psychological motivation behind political fake news sharing on Twitter. American Political Science Review 115, 3 (2021), 999–1015.
- Papakyriakopoulos and Goodman (2022) Orestis Papakyriakopoulos and Ellen Goodman. 2022. The impact of Twitter labels on misinformation spread and user engagement: Lessons from Trump’s election tweets. In WWW.
- Pierri et al. (2022) Francesco Pierri, Brea L. Perry, Matthew R. DeVerna, Kai-Cheng Yang, Alessandro Flammini, Filippo Menczer, and John Bryden. 2022. Online misinformation is linked to early COVID-19 vaccination hesitancy and refusal. Scientific Reports 12, 1 (2022), 5966.
- Plutchik (1991) Robert Plutchik. 1991. The emotions. University Press of America.
- Prochaska et al. (2023) Stephen Prochaska, Kayla Duskin, Zarine Kharazian, Carly Minow, Stephanie Blucker, Sylvie Venuto, Jevin D. West, and Kate Starbird. 2023. Mobilizing manufactured reality: How participatory disinformation shaped deep stories to catalyze action during the 2020 U.S. presidential election. In CSCW.
- Pröllochs et al. (2021a) Nicolas Pröllochs, Dominik Bär, and Stefan Feuerriegel. 2021a. Emotions explain differences in the diffusion of true vs. false social media rumors. Scientific Reports 11, 1 (2021), 22721.
- Pröllochs et al. (2021b) Nicolas Pröllochs, Dominik Bär, and Stefan Feuerriegel. 2021b. Emotions in online rumor diffusion. EPJ Data Science 10, 1 (2021), 51.
- Rao et al. (2022) Ashwin Rao, Fred Morstatter, and Kristina Lerman. 2022. Partisan asymmetries in exposure to misinformation. Scientific Reports 12, 1 (2022), 15671.
- Rathje et al. (2021) Steve Rathje, Jay J. Van Bavel, and Sander van der Linden. 2021. Out-group animosity drives engagement on social media. Proceedings of the National Academy of Sciences 118, 26 (2021), e2024292118.
- Recuero et al. (2020) Raquel Recuero, Felipe Bonow Soares, and Anatoliy Gruzd. 2020. Hyperpartisanship, disinformation and political conversations on Twitter: The Brazilian presidential election of 2018. In CSCW.
- Robertson et al. (2023b) Claire E. Robertson, Nicolas Pröllochs, Kaoru Schwarzenegger, Philip Pärnamets, Jay J. Van Bavel, and Stefan Feuerriegel. 2023b. Negativity drives online news consumption. Nature Human Behaviour 7, 5 (2023), 1–11.
- Robertson et al. (2023a) Ronald E. Robertson, Jon Green, Damian J. Ruck, Katherine Ognyanova, Christo Wilson, and David Lazer. 2023a. Users choose to engage with more partisan news than they are exposed to on Google Search. Nature (2023), 1–7.
- Roozenbeek et al. (2020) Jon Roozenbeek, Claudia R Schneider, Sarah Dryhurst, John Kerr, Alexandra LJ Freeman, Gabriel Recchia, Anne Marthe Van Der Bles, and Sander Van Der Linden. 2020. Susceptibility to misinformation about COVID-19 around the world. Royal Society open science 7, 10 (2020), 201199.
- Schmitz et al. (2023) Matheus Schmitz, Goran Muric, and Keith Burghardt. 2023. Detecting anti-vaccine users on Twitter. In ICWSM.
- Sharma et al. (2022) Karishma Sharma, Yizhou Zhang, and Yan Liu. 2022. COVID-19 vaccine misinformation campaigns and social media narratives. In ICWSM.
- Solovev and Pröllochs (2022) Kirill Solovev and Nicolas Pröllochs. 2022. Moral emotions shape the virality of COVID-19 misinformation on social media. In WWW.
- Suarez-Lledo and Alvarez-Galvez (2021) Victor Suarez-Lledo and Javier Alvarez-Galvez. 2021. Prevalence of health misinformation on social media: Systematic review. Journal of Medical Internet Research 23, 1 (2021), e17187.
- Tang et al. (2023) Chuang Tang, Shaobo (Kevin) Li, Yi Ding, Ram D. Gopal, and Guanglei Zhang. 2023. Racial discrimination and anti-discrimination: The COVID-19 pandemic’s impact on Chinese restaurants in North America. Information Systems Research (2023).
- Teng et al. (2022) Xian Teng, Yu-Ru Lin, Wen-Ting Chung, Ang Li, and Adriana Kovashka. 2022. Characterizing user susceptibility to COVID-19 misinformation on Twitter. In ICWSM.
- Van Prooijen et al. (2018) Jan-Willem Van Prooijen, Karen M. Douglas, and Clara De Inocencio. 2018. Connecting the dots: Illusory pattern perception predicts belief in conspiracies and the supernatural. European Journal of Social Psychology 48, 3 (2018), 320–335.
- Velásquez et al. (2021) N. Velásquez, R. Leahy, N. Johnson Restrepo, Y. Lupu, R. Sear, N. Gabriel, O. K. Jha, B. Goldberg, and N. F. Johnson. 2021. Online hate network spreads malicious COVID-19 content outside the control of individual social media platforms. Scientific Reports 11, 1 (2021), 11549.
- Vosoughi et al. (2018) Soroush Vosoughi, Deb Roy, and Sinan Aral. 2018. The spread of true and false news online. Science 359, 6380 (2018), 1146–1151.
- Wang and Rzeszotarski (2023) Luping Wang and Jeffrey M. Rzeszotarski. 2023. Understanding motivational factors in social media news sharing decisions. In CSCW.
- Wang et al. (2022) Xiangyu Wang, Min Zhang, Weiguo Fan, and Kang Zhao. 2022. Understanding the spread of COVID-19 misinformation on social media: The effects of topics and a political leader’s nudge. Journal of the Association for Information Science and Technology 73, 5 (2022), 726–737.
- Wang et al. (2023a) Yuping Wang, Chen Ling, and Gianluca Stringhini. 2023a. Understanding the use of images to spread COVID-19 misinformation on Twitter. In CSCW.
- Wang et al. (2023b) Ying Wang, Hongwei Zhang, Wang Gao, and Cai Yang. 2023b. Spillover effects from news to travel and leisure stocks during the COVID-19 pandemic: Evidence from the time and frequency domains. Tourism Economics 29, 2 (2023), 460–487.
- Yan et al. (2022) Muheng Yan, Yu-Ru Lin, and Wen-Ting Chung. 2022. Are mutated misinformation more contagious? A case study of COVID-19 misinformation on Twitter. In WebSci.
Supplementary Materials
Appendix S1 Cross-Correlations & Variance Inflation Factors
We additionally examine the cross-correlations and variance inflation factors among the independent variables to mitigate the concern about multicollinearity. Table S1 reports the cross-correlations among the variables in news items. Table S2 reports the cross-correlations among the variables in posts. The cross-correlations in in news items and posts are fairly small. Additionally, the variance inflation factors (VIFs) are reported in Table S3. We find that the variance inflation factors for the independent variables are all close to one and well below the critical threshold of four, which indicates that multicollinearity is not an issue of our analysis.
Falsehood | Conspiracy | Anger | Disgust | Fear | Joy | Sadness | Surprise | Words | |
Falsehood | |||||||||
Conspiracy | |||||||||
Anger | |||||||||
Disgust | |||||||||
Fear | |||||||||
Joy | |||||||||
Sadness | |||||||||
Surprise | |||||||||
Words |
Falsehood | Conspiracy | Anger | Disgust | Fear | Joy | Sadness | Surprise | Words | |
Falsehood | |||||||||
Conspiracy | |||||||||
Anger | |||||||||
Disgust | |||||||||
Fear | |||||||||
Joy | |||||||||
Sadness | |||||||||
Surprise | |||||||||
Words | |||||||||
Media | |||||||||
Verified | |||||||||
AccountAge | |||||||||
Followers | |||||||||
Followees |
Continued | Media | Verified | AccountAge | Followers | Followees |
Falsehood | |||||
Conspiracy | |||||
Anger | |||||
Disgust | |||||
Fear | |||||
Joy | |||||
Sadness | |||||
Surprise | |||||
Words | |||||
Media | |||||
Verified | |||||
AccountAge | |||||
Followers | |||||
Followees |
News | Post | |
Falsehood | ||
Conspiracy | ||
Anger | ||
Disgust | ||
Fear | ||
Joy | ||
Sadness | ||
Surprise | ||
Words | ||
Media | ||
Verified | ||
AccountAge | ||
Followers | ||
Followees | ||
Mean VIF |
Appendix S2 Robustness Check at Theme Level
We recalculate topic diversity at the theme level and repeat our analysis. Table S4 reports the estimation results for the topic diversity. Table S5 reports the estimation results for the engagement at news sharer level. Table S6 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
(1) | (2) | (3) | |
News | Post | News Post | |
Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | |
Verified | ∗∗∗ | ||
AccountAge | ∗ | ||
Followers | ∗∗∗ | ||
Followees | ∗∗∗ | ∗∗∗ | |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✗ | ✓ | ✓ |
0.147 | ✗ | ✗ |
(1) | (2) | |
Lifetime | Post Count | |
Falsehood | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗ | |
Diversity | ||
Falsehood Diversity | ||
Words | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ |
Language | ✓ | ✓ |
MonthYear | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ |
0.480 | 0.188 |
(1) | (2) | (3) | |
Repost Count | Like Count | Reply Count | |
Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Falsehood Diversity | ∗∗∗ | ∗∗ | ∗ |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✓ | ✓ | ✓ |
Appendix S3 Robustness Check with Propensity Score Matching
To reduce potential unobserved confounding factors and ensure the robustness of our results, we further conduct propensity score matching for true and false news claims. Specifically, we match true and false news claims based on the variables of emotions and the number of words, with the caliper of 0.3, “common support”, and “noreplacement” settings. Consequently, we discard 30 false news claims and achieve all biases lower than 0.05. Subsequently, we repeat our analysis based on the match true and false news claims and their associated posts on X. Table S7 reports the estimation results for the topic diversity. Table S8 reports the estimation results for the engagement at news sharer level. Table S9 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
(1) | (2) | (3) | |
News | Post | News Post | |
Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | |
Verified | ∗∗∗ | ||
AccountAge | |||
Followers | ∗∗∗ | ||
Followees | ∗∗∗ | ∗∗∗ | |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✗ | ✓ | ✓ |
0.181 | ✗ | ✗ |
(1) | (2) | |
Lifetime | Post Count | |
Falsehood | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗ | |
Diversity | ||
Falsehood Diversity | ||
Words | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ |
Language | ✓ | ✓ |
MonthYear | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ |
0.482 | 0.188 |
(1) | (2) | (3) | |
Repost Count | Like Count | Reply Count | |
Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
AccountAge | ∗∗∗ | ∗∗∗ | ∗∗ |
Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✓ | ✓ | ✓ |
Appendix S4 Robustness Check with BERTopic
To ensure the generalization of our results across the topic modeling methods, we use BERTopic, a popular unsupervised topic modeling tool, as an alternative method to automatically cluster topics.333https://maartengr.github.io/BERTopic/index.html The BERTopic model generates topic clusters for the whole news claims and posts. We use the topic clusters and their corresponding probabilities to calculate topic diversity again and repeat the analysis. Table S10 reports the estimation results for the topic diversity. Table S11 reports the estimation results for the engagement at news sharer level. Table S12 reports the estimation results for the engagement at post viewer level. All the results are robust and consistently support our main findings.
(1) | (2) | (3) | |
News | Post | News Post | |
Falsehood | ∗∗ | ||
Conspiracy | ∗ | ∗ | |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗ | ||
Verified | |||
AccountAge | |||
Followers | ∗ | ∗ | |
Followees | ∗∗∗ | ∗∗∗ | |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ||
News-level REs | ✗ | ✓ | ✓ |
0.041 | ✗ | ✗ |
(1) | (2) | |
Lifetime | Post Count | |
Falsehood | ∗∗∗ | ∗∗∗ |
Conspiracy | ∗ | |
Diversity | ∗∗∗ | ∗∗∗ |
Falsehood Diversity | ∗∗∗ | |
Words | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ |
Language | ✓ | ✓ |
MonthYear | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ |
0.508 | 0.192 |
(1) | (2) | (3) | |
Repost Count | Like Count | Reply Count | |
Falsehood | ∗ | ∗∗∗ | |
Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Falsehood Diversity | ∗ | ∗∗∗ | ∗∗∗ |
Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
Emotions | ✓ | ✓ | ✓ |
Language | ✓ | ✓ | ✓ |
MonthYear | ✓ | ✓ | ✓ |
Intercept | ∗∗∗ | ∗∗∗ | |
News-level REs | ✓ | ✓ | ✓ |

### v3

@@
-From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?
+From News Source Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis
 Abstract.
-Engagement with misinformation on social media poses unprecedented threats to societal well-being, particularly during health crises when susceptibility to misinformation is heightened in a multi-topic context.
-This paper focuses on the COVID-19 pandemic and addresses a critical gap in understanding online engagement with multi-topic misinformation at two user levels: news sharers who share source news items on social media and post viewers who engage with online news posts.
-To this end, we conduct a comprehensive analysis of fact-checked source news claims related to COVID-19 and their associated posts on X, through the lens of topic diversity and conspiracy theories.
-We find that false news, particularly when accompanied by conspiracy theories, exhibits higher topic diversity than true news.
-At the news sharer level, false news has a longer lifetime and receives more posts on X than true news.
-Additionally, the integration of conspiracy theories is significantly associated with a longer lifetime for COVID-19 misinformation.
-However, topic diversity has no significant association with news sharer engagement in terms of news lifetime and the number of posts.
-At the post viewer level, contrary to the news sharer level, news posts characterized by heightened topic diversity receive more reposts, likes, and replies.
-Notably, post viewers tend to engage more with misinformation containing conspiracy narratives: false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
-Our findings suggest that news sharers and post viewers exhibit different engagement patterns on social media regarding topic diversity and conspiracy theories, offering valuable insights into designing targeted misinformation intervention strategies at both user levels.
+Online engagement with misinformation threatens societal well-being, particularly during health crises when susceptibility to misinformation is heightened in a multi-topic context.
+Here, we focus on the COVID-19 pandemic and address a critical gap in understanding engagement with multi-topic misinformation on social media at two user levels: news source sharers (who post news items) and post viewers (who engage with news posts).
+To this end, we analyze fact-checked source news items and their associated posts on X through the lens of topic diversity and conspiracy theories.
+We find that false news, especially those containing conspiracy theories, exhibits higher topic diversity than true news.
+At news source sharer level, false news has a longer lifetime and receives more posts on X than true news, with conspiracy theories further extending its longevity.
+However, topic diversity does not significantly influence news source sharers’ engagement.
+At post viewer level, contrary to news source sharer level, posts characterized by heightened topic diversity receive more reposts, likes, and replies.
+Notably, post viewers tend to engage more with misinformation containing conspiracy narratives: false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to those without conspiracy theories.
+Our findings suggest that news source sharers and post viewers exhibit distinct engagement patterns on X, offering valuable insights into refining misinformation interventions at these two user levels.
 1. Introduction
-The spread of misinformation on social media platforms has emerged as one of the most significant global risks in recent years, posing far-reaching consequences for societies and individuals (Ecker et al., 2022).
+The spread of misinformation – i.
+e., general false or misleading information created either intentionally or unintentionally – on social media platforms has emerged as one of the most significant global risks in recent years, posing far-reaching consequences for societies and individuals (Ecker et al., 2022; Chuai et al., 2024b).
 Health misinformation, compared to misinformation in other domains, carries unique dangers due to its potential to cause harm to human bodies and lives (Chen et al., 2022).
-This issue is particularly prevalent during public health crises, such as the COVID-19 pandemic, when heightened emotions like fear and anxiety, uncertainty, and increased information consumption create fertile ground for the dissemination of false or misleading information (Chuai and Zhao, 2022; Solovev and Pröllochs, 2022; Chuai et al., 2022; Chen et al., 2023; Freiling et al., 2021).
-Throughout the COVID-19 pandemic, a substantial amount of health misinformation regarding the virus, its origin, and potential treatments or prevention strategies spread widely on social media (Erokhin et al., 2022; Sharma et al., 2022; Teng et al., 2022).
-Especially, the widespread COVID-19 vaccine misinformation on social media increases beliefs in the ineffectiveness and side effects of vaccinations, contributing to individuals’ vaccination hesitancy and refusal, thereby increasing the risk of infection (Sharma et al., 2022; Pierri et al., 2022; Enders et al., 2022).
-Furthermore, during health crises, various aspects of society and individuals become interconnected.
+This issue is particularly concerning during public health crises, such as the COVID-19 pandemic, when heightened emotions, uncertainty, and increased information consumption create fertile ground for the dissemination of false or misleading information (Chuai and Zhao, 2022; Solovev and Pröllochs, 2022; Chuai et al., 2022; Chen et al., 2023; Freiling et al., 2023).
+Additionally, as a distinct type of misinformation, conspiracy theories are deliberately fabricated and attempt to explain the causes of significant public events as secret plots by powerful and malicious groups (Douglas and Sutton, 2018; Kou et al., 2017; Suarez-Lledo and Alvarez-Galvez, 2021; Hanley et al., 2023).
+Throughout the COVID-19 pandemic, misinformation and conspiracy theories regarding the virus, its origin, and potential treatments or prevention strategies have spread widely on social media, undermining pandemic control efforts (Erokhin et al., 2022; Sharma et al., 2022b; Teng et al., 2022).
+For instance, widespread anti-vaccine conspiracy theories on social media have strengthened beliefs in the ineffectiveness and side effects of vaccinations, contributing to individuals’ vaccination hesitancy and refusal, thereby increasing the risk of COVID-19 infection (Sharma et al., 2022b; Pierri et al., 2022; Enders et al., 2022).
+Notably, during health crises, various aspects of society become interconnected.
 The discussions traverse multiple domains and demonstrate strong interconnectedness, contributing to a rich and diverse informational landscape (Miani et al., 2022; Wang et al., 2022).
-However, there remains a gap in understanding online engagement with misinformation within this multi-topic context during health crises.
-A comprehensive analysis of 13.9 million COVID-19 posts on X reveals discussions spanning more than 20 topics across the societies, encompassing areas such as politics, economy, and racism (Chandrasekaran et al., 2020).
-The diversity of topics within information content (i.
+For example, a comprehensive analysis of 13.9 million COVID-19 posts on X (formerly Twitter) reveals discussions spanning more than 20 topics across society, encompassing areas such as politics, economy, and racism (Chandrasekaran et al., 2020).
+Yet, an understanding of how users engage with misinformation within the multi-topic context during health crises is still in its infancy.
+Vosoughi et al.
+(2018) examined a large-scale of true and false news posts distributed on X from 2006 to 2017, finding that false news posts are more novel and attract more replies expressing surprise compared to true news posts.
+The novelty in their study was measured by the topic similarity between the true (false) news post a user reposted and the previous posts to which the user had been exposed before the repost.
+However, other research shows that social media users often have specific topic preferences and tend to consume similar content, especially within groups of like-minded individuals, i.
+e., echo chambers (Cinelli et al., 2021).
+Additionally, social media recommendation algorithms frequently feed users content based on their view histories and increase echo chambers, further reinforcing users’ preferences (Cinus et al., 2022; Engel et al., 2023).
+Reconciling these seemingly conflicting findings, one possible explanation is that false news may strategically connect users’ interests with other topics to pursue (hidden) agendas.
+Particularly, in the context of health crises marked by high uncertainty, diverse topics are more interconnected within discussions around crises (Chandrasekaran et al., 2020).
+Therefore, an analysis of topic interconnectedness in true and false news items during health crises is warranted.
+Previous research has shown that the diversity of topics within information content (i.
 e., multi-topic information) plays an important role in shaping perceptions and assessments of credibility through the mechanism of associative inference (Lee et al., 2023).
@@
 Moreover, the associative inference around online misinformation can evolve within a reinforcing cycle.
-Individuals collaborate in constructing a misleading version of reality, leading to offline actions that further reinforce this manufactured reality (Prochaska et al., 2023).
-The interconnected and associative multi-topic discussions during health crises underscore the importance of exploring the role of topic diversity in engagement with misinformation on social media.
-Additionally, conspiracy theories (CTs) often emerge during health crises, attempting to explain the causes of significant public events as secret plots by powerful and malicious groups (Douglas and Sutton, 2018; Kou et al., 2017; Suarez-Lledo and Alvarez-Galvez, 2021).
-The integration of conspiracy theories within misinformation narratives serves as an effective strategy to deliberately interconnect various topics (Miani et al., 2022; Hanley et al., 2023).
-This deliberate approach contributes to an increase in individuals’ beliefs in misinformation by fostering illusory pattern perception, and promoting associative inference (Van Prooijen et al., 2018; Lee et al., 2023).
+Individuals collaboratively construct a misleading version of reality and amplify alleged evidence.
+This facilitates subsequent actions that, in turn, further reinforce the manufactured reality (Prochaska et al., 2023).
+Therefore, the interconnected and associative multi-topic discussions during health crises underscore the potential role of topic diversity in driving engagement with misinformation on social media.
+Based on this rationale, we address the following research questions to explore topic diversity in misinformation (RQ1.1) and examine whether it shapes user engagement (RQ1.2):
+RQ1.1: Does misinformation exhibit higher topic diversity compared to true information?
+RQ1.2: Is topic diversity positively associated with user engagement, and is this association stronger for misinformation than for true information?
+Moreover, the integration of conspiracy theories within misinformation narratives serves as an effective strategy to deliberately interconnect various topics (Miani et al., 2022; Hanley et al., 2023).
+This deliberate approach contributes to an increase in individuals’ beliefs in misinformation by fostering illusory pattern perception and promoting associative inference (Van Prooijen et al., 2018; Lee et al., 2023).
 In the early months of the COVID-19 pandemic, social media platforms were inundated with various conspiracy theories related to the origins, spread, and treatment of the virus (Erokhin et al., 2022).
-The propagation of misinformation with coordinated conspiracies promotes distrust in public health authorities and skepticism about the safety and effectiveness of COVID-19 vaccines (Sharma et al., 2022; Schmitz et al., 2023).
+The propagation of misinformation with coordinated conspiracies has promoted distrust in public health authorities and skepticism about the safety and effectiveness of COVID-19 vaccines (Sharma et al., 2022b; Schmitz et al., 2023).
 While the proliferation of conspiracy theories on social media has raised significant concerns, little is known about how the incorporation of conspiracy theories in misleading narratives differentiates individuals’ engagement on social media compared to other types of misinformation (Kim and Kim, 2023).
-This gap calls for exploring how conspiracy theories affect engagement with misinformation during health crises.
-Research questions: In this paper, we aim to provide a comprehensive understanding of online engagement with multi-topic misinformation during the early stages of the COVID-19 pandemic, focusing on topic diversity and conspiracy theories.
-To this end, we investigate the following Research Questions (RQs):
-RQ1: Does misinformation exhibit higher topic diversity compared to true information?
-RQ2: Does misinformation containing conspiracy theories exhibit higher topic diversity compared to misinformation without conspiracy theories?
-RQ3: Does misinformation with higher topic diversity receive more engagement compared to true information?
-RQ4: Does misinformation containing conspiracy theories receive more engagement compared to misinformation without conspiracy theories?
-Of note, we examine engagement at two levels: news sharer level and post viewer level.
-News sharers create posts on X that link to source news items outside of the platform.
-Post viewers engage with news posts within the platform through the reactions of reposting, liking, or replying.
+Given this, we further investigate the following research questions to analyze whether misinformation containing conspiracy theories exhibits higher topic diversity (RQ2.1) and receives more engagement (RQ2.2) compared to misinformation without conspiracy theories:
+RQ2.1: Does misinformation containing conspiracy theories exhibit higher topic diversity compared to misinformation without conspiracy theories?
+RQ2.2: Does misinformation containing conspiracy theories receive more engagement compared to misinformation without conspiracy theories?
+To address our research questions, we perform a comprehensive analysis of online engagement with multi-topic misinformation during the early stages of the COVID-19 pandemic, focusing on topic diversity and conspiracy theories.
+It is worth noting that we examine engagement at two levels: news source sharer level and post viewer level.
+News source sharers create posts on X that link to source news items, while post viewers encounter these news posts within the platform and then engage by reposting, liking, or replying.
 The two user groups may have different motivations for engagement (Hanley et al., 2023; Teng et al., 2022; Ballard et al., 2022).
-Additionally, while previous studies have extensively explored post viewer engagement (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021a; Solovev and Pröllochs, 2022), research on news sharer engagement remains limited.
+Additionally, while previous studies have extensively explored post viewer engagement on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021a; Solovev and Pröllochs, 2022), research on news source sharer engagement remains limited.
 Understanding this under-investigated area is crucial to complement our knowledge of news-sharing behavior.
@@
 These source news claims span six languages and cover the initial period of COVID-19 from January to September 2020.
-At news sharer level, the engagement metrics include the number of posts linking to a specific news item and the lifetime of the news item, which indicates the duration from the creation of the first post to the latest post.
+At news source sharer level, the engagement metrics include the number of posts linking to a specific news item and the lifetime of the news item from the creation of its first post to its latest post.
 At post viewer level, the engagement metrics include repost count, like count, and reply count.
@@
 Contributions: Our work reveals that elevated topic diversity has been a characteristic of COVID-19 misinformation, particularly for those accompanied by conspiracy narratives.
-This characteristic becomes more significant within the false news posts on X compared to their source news claims.
-At post viewer level, the heightened topic diversity and the integration of conspiracy theories significantly correlate with increased engagement (i.
+This characteristic becomes more significant within false news posts on X compared to their source news claims.
+At post viewer level, heightened topic diversity and the integration of conspiracy theories significantly correlate with increased engagement (i.
 e., reposts, likes, and replies) with misinformation on X.
 For instance, false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
-However, at news sharer level, topic diversity and conspiracy theories have no significant association with the post count of source news items.
-This difference suggests that news sharers and post viewers may have different engagement patterns.
-Our findings offer valuable insights into understanding the engagement with multi-topic misinformation from news sharers to post viewers on social media during health crises.
-Additionally, by highlighting the different engagement patterns at news sharer and post viewer levels in terms of topic diversity and conspiracy theories, this research informs the development of targeted interventions and strategies at both user levels to mitigate engagement with false information and foster a more resilient society.
+However, at news source sharer level, topic diversity and conspiracy theories have no significant associations with the number of posts that source news items receive.
+This difference suggests that news source sharers and post viewers have distinct engagement patterns.
+Our findings offer valuable insights into understanding engagement with multi-topic misinformation from news source sharers to post viewers on social media during health crises.
+Additionally, by highlighting the different engagement patterns at news source sharer and post viewer levels in terms of topic diversity and conspiracy theories, this research informs the development of targeted interventions and strategies at both user levels to mitigate engagement with false information and foster a more resilient society.
 2. Background and Related Work
 2.1. Misinformation on Social Media
-Social media platforms have transformed information exchange, enabling public users to create and engage with online content globally and instantaneously, albeit without the quality control mechanisms inherent in traditional media.
-This has led to a significant rise in misinformation.
-Particularly, the prevalence of misinformation on social media has posed far-reaching consequences across many critical societal domains ranging from politics to public health, becoming a concerning global issue in recent years (Ecker et al., 2022; Pierri et al., 2022; Enders et al., 2022; Green et al., 2022).
-Consequently, social media platforms are under increasing pressure to implement measures to reduce the spread of misinformation.
-Understanding how users engage with misinformation and contribute to its viral spread on social media can help social media platforms to design targeted interventions.
+Social media platforms have transformed information exchange, enabling public users to create and engage with online content globally and instantaneously.
+However, the absence of quality control mechanisms inherent in traditional media has led to a significant rise in misinformation on social media.
+Particularly, the spread of misinformation on social media has posed far-reaching consequences across many critical societal domains ranging from politics to public health, thus becoming a concerning global issue in recent years (Ecker et al., 2022; Pierri et al., 2022; Enders et al., 2022; Green et al., 2022).
+Consequently, social media platforms are under increasing pressure to implement effective measures to reduce the spread of misinformation.
+Understanding how users engage with misinformation and contribute to its viral spread on social media can help social media platforms design targeted interventions.
 Prior research has shown that online misinformation spreads farther, faster, deeper, and more broadly compared to true information, as evidenced by the examination of online posts fact-checked by third-party organizations (Vosoughi et al., 2018).
@@
 Several works have focused on understanding what characterizes misinformation and contributes to its viral spread.
-For instance, misinformation is more likely to incorporate negative emotions, such as anger, which have been identified as a main driver for information diffusion (Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b).
-Additionally, there is a partisan asymmetry that misinformation tends to be right-leaning and promote conservative positions (Chuai et al., 2023).
+For instance, misinformation is more likely to incorporate negative emotions, such as anger, which has been identified as an important driver for information diffusion (Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b).
+Additionally, there is a partisan asymmetry that misinformation tends to be right-leaning and promote conservative positions (Chuai et al., 2025).
 Social media users are more likely to believe and share information that is congruent to their political ideologies (Garrett and Bond, 2021; Rao et al., 2022; Robertson et al., 2023a; Osmundsen et al., 2021; Jenke, 2023).
 Therefore, right-leaning users are more likely to share misinformation on social media, compared to left-leaning users (Nikolov et al., 2021).
-More importantly, uncertain situations, such as emergency public events or crises, create fertile ground for online misinformation to thrive.
-During such times, people are eager to know more information about the situations and thus more likely to believe and share related misinformation, which can destroy the trust in the control policies and cause detrimental harm to societies and individuals.
+More importantly, uncertain situations, such as public emergencies or crises, create fertile ground for online misinformation to thrive.
+During such times, people are eager to know more information about the situation and are thus more likely to believe and share related misinformation, which can undermine trust in control policies and cause detrimental harm to societies and individuals.
 2.2. Misinformation During Public Events/Crises
@@
 However, democracy can be undermined by the widespread dissemination of dedicated online misinformation aimed at pursuing specific political agendas.
-It has been found that people are exposed to misinformation more often during elections than usual (Allcott and Gentzkow, 2017).
+It has been found that people are exposed to misinformation more often during elections than usual (Allcott and Gentzkow, 2017; Chuai et al., 2025).
 Election-related misinformation is often polarized, which can increase people’s susceptibility and affect their voting behaviors (Rathje et al., 2021; Nikolov et al., 2021; Recuero et al., 2020).
@@
 The spread of misinformation during health crises is particularly severe, as it can directly harm human health and even cost lives (Chen et al., 2022).
-Throughout the COVID-19 pandemic, a significant volume of misinformation regarding a wide variety of COVID-19 related topics has circulated widely on social media platforms.
-For example, the widespread COVID-19 vaccine misinformation reduces users’ beliefs in its effectiveness, impeding the fight against the spread of COVID-19 (Enders et al., 2022; Loomba et al., 2021; Pierri et al., 2022).
+Throughout the COVID-19 pandemic, misinformation on a wide variety of COVID-19 related topics has circulated widely on social media platforms.
+For example, the widespread misinformation on the COVID-19 vaccine reduces users’ beliefs in its effectiveness, impeding the fight against the spread of COVID-19 (Enders et al., 2022; Loomba et al., 2021; Pierri et al., 2022).
 A global survey has revealed that a substantial part of people consider COVID-19 misinformation as highly reliable (Roozenbeek et al., 2020).
-Understanding how misinformation spreads and engages users during such public events and crises is crucial for developing effective intervention strategies for emergency management.
+Understanding how misinformation spreads and attracts users during such public events and crises is crucial for developing effective intervention strategies for emergency management.
 2.3. Engagement With COVID-19 Misinformation on Social Media
-The engagement with misinformation during health crises, particularly COVID-19, posed significant threats to the pandemic control and public safety.
+The engagement with misinformation during health crises, particularly COVID-19, poses significant threats to crisis control and public safety.
 Several studies have scrutinized the factors that can influence engagement with COVID-19 misinformation on social media.
-Recognizing emotions as a significant driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a), research in this body has predominantly focused on the role of emotions during COVID-19 (Solovev and Pröllochs, 2022; Chuai et al., 2022; Freiling et al., 2021; Charquero-Ballester et al., 2021).
+Recognizing emotions as a significant driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a), research in this body has predominantly focused on the role of emotions during COVID-19 (Solovev and Pröllochs, 2022; Chuai et al., 2022; Freiling et al., 2023; Charquero-Ballester et al., 2021).
 For example, Solovev and Pröllochs (2022) conducted a large-scale computational analysis on the role of moral emotions in the spread of COVID-19 misinformation.
@@
 (2022) focused on the specific topics in the engagement with COVID-19 vaccine misinformation.
-They found that safety concern was the most significant content theme but negatively associated with likes and shares.
-However, during COVID-19, various aspects of society and individuals became interconnected.
+They found that safety concerns are the most significant content theme but are negatively associated with likes and shares.
+However, during COVID-19, various aspects of society became interconnected.
 COVID-19 misinformation covered a broad spectrum of topics, including public health, politics, medicine, vaccines, and so forth (Wang et al., 2022; Chandrasekaran et al., 2020).
@@
 2.4. Multi-Topic Context and Conspiracy Theories During COVID-19
-The discussion on COVID-19 during the pandemic has not only focused on the direct health impacts but has also spilled over into various social, economic, and political domains.
-Within the social domain, at the beginning of the pandemic, rumors and conspiracy theories about the origin of the virus in China were rampant on social media, which exploits anti-Asian racism and hate speech on social media (Huang et al., 2023).
-Using COVID-19 as a weapon, malicious misinformation fosters online hate to spread quickly beyond the control of social media platforms (Velásquez et al., 2021).
-The rising anti-Asian racism during the COVID-19 pandemic knocked Asian businesses (Huang et al., 2023; Luca et al., 2024; Tang et al., 2023).
-In the economic domain, the pandemic hit the global economy, particularly impacting sectors such as travel and leisure (Wang et al., 2023b).
-Additionally, the scientific uncertainty surrounding the COVID-19 pandemic fuels the weaponization of COVID-19 for politicization (Kreps and Kriner, 2020).
+The discussion on COVID-19 during the pandemic not only focused on the direct health impacts but also spilled over into various social, economic, and political domains.
+Within the social domain, at the beginning of the pandemic, rumors and conspiracy theories about the origin of the virus in China were rampant on social media, exploiting anti-Asian racism and hate speech on social media (Huang et al., 2023).
+Using COVID-19 as a weapon, malicious misinformation fostered online hate to spread quickly beyond the control of social media platforms (Velásquez et al., 2021).
+The rising anti-Asian racism during the COVID-19 pandemic knocked Asian businesses (Huang et al., 2023; Luca et al., 2024; Tang et al., 2024).
+In the economic domain, the pandemic hit the global economy, particularly impacting sectors such as travel and leisure (Wang et al., 2023).
+Additionally, the scientific uncertainty surrounding the COVID-19 pandemic fueled the weaponization of COVID-19 for politicization (Kreps and Kriner, 2020).
 Some individuals and groups have sought to exploit the pandemic to spread extremist ideologies and radicalize others (Davies et al., 2023).
-The high degree of politicization within COVID-19 information contributes to polarization in attitudes towards the pandemic and even to political ideologies (Hart et al., 2020).
-For example, the discussion on vaccination become gradually polarized, expanding far beyond the medical discourse and entering the domain of organizational politics (Dehghan and Nagappa, 2022).
-Taken together, the COVID-19 pandemic is incorporated into many other societal domains to pursue specific interests, yielding significant spillover effects on the stability of societies.
+The high degree of politicization within COVID-19 information contributed to polarization in attitudes towards the pandemic and even to political ideologies (Hart et al., 2020).
+For instance, the debate over vaccination became increasingly polarized during the pandemic, expanding far beyond the medical discourse and entering the domain of organizational politics (Dehghan and Nagappa, 2022).
+Taken together, the COVID-19 pandemic has been incorporated into many other societal domains to pursue specific interests, yielding significant spillover effects on the stability of societies.
 In this paper, we use topic diversity to denote the extent to which the information involves multiple topics and explore its impact on the engagement with misinformation on social media during COVID-19.
@@
 This repository provides an extensive COVID-19 dataset that consists of source news claims and their corresponding engagement on X.
-Initially, the repository gathers these news claims and their veracity verdicts from reputable third-party fact-checking organizations worldwide, including Snopes and Poynter, as well as official health websites.
-To keep the quantity of the source claims in each language, the repository filters six languages: English, Spanish, Portuguese, Hindi, French, and Italian, which enables a broad and diverse examination of misinformation across linguistic and cultural contexts.
+Initially, the repository gathers these news claims and their veracity verdicts from popular third-party fact-checking organizations worldwide, including Snopes and Poynter, as well as official health websites.
+To keep the quantity of the source claims in each language, the repository filters six languages: English, Spanish, Portuguese, Hindi, French, and Italian.
+This enables a broad and diverse examination of misinformation across linguistic and cultural contexts.
 Subsequently, online posts that explicitly reference these news sources are collected from X through its advanced search API.
-Notably, the news sources in the dataset are linked to the original news URLs.
-However, due to the diverse range of news items collected from various languages and web pages, precisely parsing the original news pages and obtaining their content is challenging.
+Notably, while the news sources in the dataset are linked to the original news URLs, precisely parsing the original news pages and obtaining their content is challenging due to the diverse range of news items collected from various languages and web pages.
 Instead, the dataset provides news claims, representing the main opinions and topics.
-Data preprocessing: We begin by translating all the non-English texts in source news claims and their associated posts into English using the Googletrans Python package.
+Data preprocessing: We begin by translating all the non-English texts in source news claims and their associated posts into English using Google Translate API via Googletrans.111Googletans library is available at https://github.com/ssut/py-googletrans.
+It is a free and unlimited Python library that implements Google Translate API and uses the same servers that translate.google.com uses.
+Google Translate has a high performance in translating between English and non-English languages, retaining overall meaning in 82.5% of the translations (Taira et al., 2021; Sebo and de Lucia, 2024).
 Subsequently, we perform text preprocessing by removing non-alphabetic and non-numerical characters, stop words, and URLs.
@@
 To ensure a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we utilize these keywords as seeds to identify similar words and construct a topic lexicon based on a Word2Vec model.
-Notably, after reexamination, we remove the topic about locations that have no specific meanings and consider the remaining 25 topics across 10 themes (Table 1).
+Upon reevaluating the associations between topic categories and their themes, we find that the keywords in the topic “Hotspots and location” relate solely to places, such as “cities” and “location,” not adequately reflecting the theme “Spread and growth” in terms of transmission modes or case numbers.
+Therefore, in the main analysis, we remove the topic “Hotspots and location” and focus on the remaining 25 topics across 10 themes (Table 1).222In a subsequent robustness check, we conduct the analysis again using all the original topics and keywords.
+The results are robust and consistently support our findings.
+The full estimation results are reported in Suppl.
+S3.
 | Theme | Topic |
@@
 | testing | |
-Impact on health care sector | Impact on hospitals and clinics, policy changes, |
+| Impact on health care sector | Impact on hospitals and clinics, health policy, |
 | frontline workers | |
@@
 | employment, impact on business | |
-Political impact | Political impact |
+| Political impacts | Political impacts |
 | Racism | Racism |
-Word2Vec model: To train the Word2Vec model, we additionally collect a large-scale COVID-19 dataset from a public repository on Kaggle.111https://www.kaggle.com/datasets/smid80/coronavirus-covid19-tweets This dataset contains posts in multiple languages, posted from March 9 to April 30, 2020 on X.
-Given the size of the dataset, we only filter English posts and do not involve the translation process.
-By combining this dataset with MM-COVID, we obtain a total of English posts.
-After removing non-alphabetic and non-numerical characters as well as URLs in the posts, we utilize the refined corpus to train a Word2vec model with a vector size of 200 using the gensim library.222The model is trained on the High Performance Computing platform at the University of Luxembourg.
-The examples of the keywords and their ten most similar words are shown in Table 2.
 | bioweapon | vaccine | election | quarantine |
@@
 | biowar | antibodies | generalelection | isolating |
-Topic lexicon: We expand our set of topic keywords by incorporating words that exhibit similarities of more than 0.6 with the initial topic seeds.
-This process yields 828 additional topic keywords.
-Next, we conduct a manual review of these keywords, identifying that 90.6% of them are indeed relevant to the corresponding topic seeds and effectively represent the associated topics.
-Consequently, we integrate the topic seeds with these related keywords to create a comprehensive topic lexicon.
+Word2Vec model: To train the Word2Vec model, we additionally collect a large-scale COVID-19 dataset from a public repository on Kaggle.333The dataset is available at https://www.kaggle.com/datasets/smid80/coronavirus-covid19-tweets.
+This dataset contains posts in multiple languages, posted from March 9 to April 30, 2020 on X.
+Given the size of the dataset, we only filter English posts and do not involve the translation process.
+By combining this dataset with MM-COVID, we obtain a total of English posts.
+After removing non-alphabetic and non-numerical characters as well as URLs in the posts, we utilize the refined corpus to train a Word2vec model with a vector size of 200 using the gensim library.444The model is trained on the High Performance Computing platform at the University of Luxembourg.
+The examples of the keywords and their ten most similar words are shown in Table 2.
+Topic lexicon: We expand our set of topic keywords by incorporating words that exhibit cosine similarities of more than 0.6 with at least one of the initial topic seeds.
+This process yields 828 additional topic keywords, and they are automatically categorized into the same topics as their corresponding topic seeds.
+Additionally, we conduct a manual review of these keywords and remove ambiguous terms and abbreviations.
+For example, we remove the keyword “reason” from the topic “Alternative causes,” as it does not specify a particular alternative cause for COVID-19, whereas the keyword “chinavirus” is an effective term for this topic category.
+Similarly, we remove ambiguous abbreviations such as “pres” in the topic “Political impacts.” Ultimately, we retain 90.6% of the 828 keywords and integrate them with the topic seeds to create an extensive topic lexicon.555We repeat our analysis using the keywords without manual validation.
+The results remain consistent, further enhancing the robustness of our findings.
+The full estimation results are reported in Suppl.
+S3.
 Topic identification: We utilize the topic lexicon to identify topics discussed in both news claims and their associated posts.
@@
 Compared to true news claims, false news claims contain 56.1% more co-occurrence edges and are more interconnected, indicating higher interconnectedness and suggesting greater topic diversity in false news claims.
+For example, the topic “Political impacts” is connected to 5 topics in true news claims, whereas it connects to 10 additional topics, such as “Death reports” and “Drugs and vaccines,” in false news claims (Fig.
+1(c)).
 Building upon previous research that used the Gini coefficient to measure the inequality of topic distribution in the content (Miani et al., 2022), we quantitatively measure topic diversity () as:
@@
 3.4. COVID-19 Conspiracy Theories
-Conspiracy theories have been often thought of as a subset of misinformation (Hanley et al., 2023).
-Following the method commonly adopted by previous related research (Erokhin et al., 2022; Moffitt et al., 2021; Sharma et al., 2022), we select a set of keywords (phrases) from the topic lexicon to identify the conspiracy theories embedded in the false news items and associated posts.
-Specifically, we consider eight COVID-19 conspiracy theory categories that connect to at least one topic in our analysis: the sources and origins of the virus that caused the outbreak of COVID-19 (VirusOrigin), the connection between 5G and COVID-19 (5G), the role of big pharmaceutical companies (BigPharma) and Bill Gates (BillGates) in the spread of COVID-19, the scope of the pandemic is exaggerated (Exaggeration), the claim that hospitals were actually empty (FilmYourHospital), the idea that the genetically modified crops led to the emergence of COVID-19 (GMO), and skepticism regarding the ineffectiveness and side-effects of vaccines (Vaccines).
-As shown in Table 3, we match the keywords with the lemmatized words in each news claim or post, and for phrases, we directly match them with the original translated text.
+The outbreak of COVID-19 has fueled widespread conspiracy theories with misleading claims about its origin, treatment, and other related topics.
+We summarize the conspiracy theories identified in previous research (Erokhin et al., 2022; Moffitt et al., 2021; Sharma et al., 2022b) and use a set of keywords and phrases to identify conspiracy theories embedded in false news claims and subsequent posts.
+Specifically, we consider eight COVID-19 conspiracy theory categories that connect to at least one topic in our analysis: the sources and origins of the virus that caused the outbreak of COVID-19 (VirusOrigin), the connection between 5G and COVID-19 (5G), the role of big pharmaceutical companies (BigPharma) and Bill Gates (BillGates) in the spread of COVID-19, the claim that the scope of the pandemic was exaggerated (Exaggeration), the claim that hospitals were actually empty (FilmYourHospital), the idea that the genetically modified crops led to the emergence of COVID-19 (GMO), and the skepticism regarding the ineffectiveness and side-effects of vaccines (Vaccines).
+We combine the conspiracy-specific terms adopted by prior studies (Erokhin et al., 2022; Moffitt et al., 2021) and compile an extensive list of keywords and phrases to identify the eight conspiracy theories (Table 3).
+Subsequently, we match these keywords with the lemmatized words in each news claim or post, and for phrases, we directly match them with the original translated text.
 We find that false news claims and false news posts incorporate keywords (phrases) indicating conspiracy theories.
-Given that the share of conspiracy posts ranges from 0.6% to 18% (Erokhin et al., 2022), the 6% share of posts containing conspiracy theories in our dataset is reasonable.
-Upon a manual check of news items that contain keywords or phrases related to conspiracy theories, we find that the accuracy is 92.1%, which suggests that the selected keywords (phrases) can effectively identify the conspiracy theories integrated into the misinformation (Erokhin et al., 2022; Moffitt et al., 2021).
+Given that the shares of conspiracy posts range from 0.6% to 18% in previous studies (Erokhin et al., 2022), the 6% share of posts containing conspiracy theories in our dataset is reasonable.
+Upon a manual check of the source news items that contain keywords or phrases related to conspiracy theories, we find that the accuracy is 92.1%.
+Therefore, the selected keywords (phrases) can effectively identify conspiracy theories in our dataset.
 | Conspiracy theory | Keywords |
@@
 | GMO | gmo, genetically modified |
-3.5. The Engagement From Source News to Posts on X
+3.5. User Engagement on X
 On the platform of X, users can interact with posts they view through reposting, liking, or replying.
-Existing research often uses the number of reposts to study engagement with misinformation on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b; Drolsbach and Pröllochs, 2023; Chuai et al., 2024).
-Additionally, the number of likes and the number of replies are also important indicators of engagement with online posts (Chuai et al., 2024; Papakyriakopoulos and Goodman, 2022).
-Notably, news items outside of X enter the social environment through news sharers on X.
-Multiple posts on X can link to the same original news items, which increases the visibility of the news items.
-However, few studies examine engagement with misinformation from the perspective of source news items to posts.
-In this paper, we consider user engagement at two levels: news sharer level and post viewer level (Fig.
+Existing research typically uses the number of reposts to study engagement with misinformation on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b; Drolsbach and Pröllochs, 2023; Chuai et al., 2024b).
+Additionally, the number of likes and the number of replies are also important indicators of engagement with online posts (Chuai et al., 2024b; Papakyriakopoulos and Goodman, 2022).
+However, few studies examine engagement with misinformation from the perspective of news source sharing.
+Multiple posts on X can link to the same source news items, which increases the visibility of the news items.
+In this paper, we consider user engagement at two levels: news source sharer level and post viewer level (Fig.
 2).
-At news sharer level, news items are shared via posts on X.
+At news source sharer level, news items are shared by news source sharers via posts on X.
 The engagement metrics include the number of posts and the lifetime of news items from the creation of the first post to the creation of the latest post.
 At post viewer level, engagement is measured by how viewers interact with these posts.
-Post viewers can repost, like, and reply to the posts that are created by news viewers and contain links to the original news items.
+Post viewers can repost, like, and reply to posts that are created by news source sharers and contain links to the original news items.
 3.6. Empirical Models
-We employ multiple explanatory regression models to explore the research questions separately. To this end, we define the following dependent variables to measure engagement:
+We employ multiple explanatory regression models to explore our research questions separately.
+To this end, we define the following variables to measure engagement at news source sharer level and post viewer level.
+The engagement at news source sharer level:
 : The duration (in hours) from the creation of the first post to the creation of the latest post on X for the specific news item.
 : A count variable indicating the number of posts associated with the specific news item.
+The engagement at post viewer level:
 : A count variable indicating the number of reposts received by the specific post.
@@
 : A count variable indicating the number of replies received by the specific post.
-The key independent variables include:
+Additionally, we define the following key variables to address our research questions:
 : A dummy variable indicating whether the news claim is false ( 1) or true ( 0).
@@
 : The topic diversity of the news claim or post.
-We also control for the variables from post and user characteristics:
+We also control for the variables from content characteristics that can affect user engagement:
 : A continuous-like count variable indicating the number of words in the text of the news claim or post.
+This control variable is commonly used to account for the content richness (Miani et al., 2022).
 : A dummy variable indicating whether the post has media elements ( 1) or not ( 0).
-: A dummy variable indicating whether the post account is verified ( 1) or not ( 0).
-: A continuous-like count variable indicating the age (in days) of the post account.
-: A continuous-like count variable indicating the number of followers of the post account.
-: A continuous-like count variable indicating the number of followees of the post account.
-Additionally, emotions are considered as a main driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a).
+Previous research has found that media elements, such as images and videos, can facilitate user engagement (Zhou et al., 2021).
+Emotions are considered to be the main driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a).
 Our analysis considers six cross-cultural emotions – anger, disgust, fear, joy, sadness, and surprise – out of the eight basic emotions (Chuai and Zhao, 2022; Plutchik, 1991).
@@
 If the news claims (posts) have no emotion words, all the weights of emotions are set to 0.
-The descriptive statistics for the variables are shown in Table 4.
+Finally, the posters on social media play an important role in driving engagement with their posts (Vosoughi et al., 2018; Chuai and Zhao, 2022).
+We use the following four variables to account for poster characteristics:
+: A dummy variable indicating whether the post account is verified ( 1) or not ( 0).
+: A continuous-like count variable indicating the age (in days) of the post account.
+: A continuous-like count variable indicating the number of followers of the post account.
+: A continuous-like count variable indicating the number of followees of the post account.
+The descriptive statistics for the above variables are shown in Table 4.
 | (1) | (2) | |
@@
 | Followees | 3,031.517 (15,099.135) | 2,361.965 (15,535.361) |
-Topic diversity (RQ1 & RQ2): We use as the dependent variable to analyze the topic diversity in true and false news from original news claims to online posts.
-In terms of the claims, the linear regression model that incorporates the main independent variables of and is specified as:
+Topic diversity (RQ1.1 & RQ2.1): We use as the dependent variable to analyze topic diversity in true and false news from original news claims to online posts.
+In terms of claims, the linear regression model that incorporates the main independent variables of and is specified as:
 where indicates six basic emotions. Additionally, denotes language-specific fixed effects, and denotes month-year fixed effects.
@@
 Therefore, we specify a liner regression model incorporating news-specific random effects:
-where indicates news-specific random effects. Additionally, we also incorporate , poster characteristics as control variables in the model.
-Engagement at news sharer level – lifetime and post count (RQ3 & RQ4): The first layer of engagement is from source news claims to news posts on X.
+where indicates news-specific random effects.
+Additionally, we incorporate , poster characteristics as control variables.
+In the above two models, the coefficient estimate of () is to address RQ1.1, and the coefficient estimate of () is to address RQ2.1.
+Engagement at news source sharer level – news lifetime and post count (RQ1.2 & RQ2.2): The first layer of engagement is from source news claims to news posts on X.
 We examine how topic diversity and conspiracy theories influence engagement with misinformation in terms of news lifetime and post count.
@@
 (4).
-Engagement at post viewer level – reposts, likes, and replies (RQ3 & RQ4): To examine the effects of topic diversity and conspiracy theories on social engagement (i.
-e., reposts, likes, and replies) with misinformation on X/Twitter, we specify a negative binomial regression model with language-specific fixed effects, month-year fixed effects, and news-specific random effects:
-where indicates all the independent variables, and signifies the dependent variables, including , and . All the continuous variables in the above models are z-standardized.
+Engagement at post viewer level – reposts, likes, and replies (RQ1.2 & RQ2.2): To examine the effects of topic diversity and conspiracy theories on social engagement (i.
+e., reposts, likes, and replies) with misinformation on X, we specify a negative binomial regression model with language-specific fixed effects, month-year fixed effects, and news-specific random effects:
+where indicates all the independent variables, and signifies the dependent variables, including , and .
+In the above models for engagement, the coefficient estimates of (), (), and () are to address RQ1.2, and the coefficient estimate of () is to address RQ2.2.
+All the continuous variables in these models are z-standardized.
 4. Empirical Results
-4.1. Topic Diversity (RQ1 & RQ2)
-Summary statistics: Fig.
-3(a) shows that false news claims (mean of 0.086) exhibit slightly higher topic diversity compared to true news claims (mean of 0.082, , ).
-Notably, the difference in topic diversity between true and false news items is magnified at the post level, where the mean topic diversity in false news posts is 0.113, compared to 0.079 in true news posts (Fig.
+4.1. Topic Diversity (RQ1.1 & RQ2.1)
+4.1.1. Descriptive statistics.
+We start by examining topic diversity from news claims to news posts (Figs.
+3(a)–3(c)).
+Fig.
+3(a) shows that false news claims (mean of 0.086) exhibit 4.9% higher topic diversity compared to true news claims (mean of 0.082, , ).
+This difference is significantly magnified (43.0% higher) at the post level, where the average topic diversity in false news posts is 0.113, compared to 0.079 in true news posts (Fig.
 3(b), , ).
-Additionally, from claims to posts (Fig.
-3(c)), topic diversity in false news items (mean of 0.031) increases more than that in true news items (mean of 0.012, , ).
+Additionally, we examine the paired changes in topic diversity from claims to their corresponding posts (Fig.
+3(c)), finding that topic diversity in false news items (increase of 0.031) increases 158.3% more than that in true news items (increase of 0.012, , ).
 For false news items containing conspiracy theories (CTs), the topic diversity at the claim level has no statistically significant difference compared to the overall false news items (Fig.
@@
 However, Fig.
-3(c) indicates a higher proportion of the segment with most significant increases in topic diversity for CT news items, compared to false news items ( , ), which may lead to the significant difference in topic diversity between CT news and general false news at the post level (Fig.
+3(c) indicates that the most significant increases in topic diversity from claims to posts are more likely to be in CT news items, compared to overall false news items ( , ).
+This may lead to the significant difference in topic diversity between CT news and general false news at the post level (Fig.
 3(b), , ).
-Regression results: In Column (1) of Table 5, the coefficient estimate of is significantly positive ( , ).
-This implies that false news exhibits higher topic diversity compared to true news at the claim level, aligning with the observations in Fig.
+4.1.2. Regression results.
+We further conduct regression analysis to estimate the effects of and on topic diversity (Figs.
+3(d)–3(f)).
+In Fig.
+3(d), the coefficient estimate of is significantly positive ( , ).
+This implies that the topic diversity in false news is, on average, 0.388 standard deviations higher than in true news at the claim level, consistent with the observations in Fig.
 3(a).
-Additionally, the coefficient estimate of is not statistically significant, which indicates that the topic diversity in false news claims containing conspiracy theories is not significantly different from that in false news claims without conspiracy theories.
+Additionally, the coefficient estimate of is not statistically significant, which indicates that the topic diversity in false news claims containing conspiracy theories is not significantly different from that in false news claims without conspiracy theories (Fig.
+3(d)).
 Subsequently, we explore the topic diversity between true and false news at the post level.
-In Column (2) of Table 5, the coefficient estimate of is significantly positive ( , ), suggesting that false news maintains higher topic diversity than true news at the post level.
-(1) | (2) | (3) | |
-News | Post | News Post | |
-Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Conspiracy | ∗∗∗ | ∗∗∗ | |
-Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Media | ∗∗∗ | ∗∗∗ | |
-Verified | ∗∗∗ | ||
-AccountAge | |||
-Followers | ∗∗∗ | ||
-Followees | ∗∗∗ | ∗∗∗ | |
-Emotions | ✓ | ✓ | ✓ |
-Language | ✓ | ✓ | ✓ |
-MonthYear | ✓ | ✓ | ✓ |
-Intercept | ∗∗∗ | ∗∗∗ | |
-News-level REs | ✗ | ✓ | ✓ |
-0.179 | ✗ | ✗ |
-Notably, as observed in Fig.
-3(b) and the coefficient estimates of in Columns (1) and (2), the difference between true and false news posts is larger than that between true and false news claims.
-Additionally, the coefficient estimate of becomes significantly positive ( , ), which suggests that false news posts containing conspiracy theories exhibit even higher topic diversity than false news posts without conspiracy theories.
+In Fig.
+3(e), the coefficient estimate of is significantly positive ( , ), suggesting that the topic diversity in false news is, on average, 0.751 standard deviations higher than in true news at the post level.
+Notably, as observed in the CCDF in Fig.
+3(b) and the coefficient estimates of in Figs.
+3(d)–3(e), the difference between true and false news posts is larger than that between true and false news claims.
+Additionally, the coefficient estimate of in Fig.
+3(e) becomes significantly positive ( , ).
+This suggests that false news posts containing conspiracy theories exhibit even higher topic diversity (0.136 standard deviations more) than false news posts without conspiracy theories.
 There are two possible reasons for the larger difference in topic diversity between true and false news at the post level compared to the claim level.
@@
 Another is that false news has a bigger increase in topic diversity from the claim level to the post level, compared to true news.
-Given the second reason, we further examine the change in topic diversity from the source news claims to the corresponding posts.
-In Column (3) of Table 5, the positive coefficient estimate of is statistically significant ( , ).
-It suggests that the increase of topic diversity in false news from the claim level to the post level is bigger than that in true news, which could partially explain the enlarged difference in topic diversity between true and false news posts, compared to their source claims.
-In addition, the coefficient estimate of in Column (3) of Table 5 is also significantly positive ( , ).
-It means that the topic diversity in false news that contains conspiracy theories increases more than that in false news without conspiracy theories from the claim level to the post level.
-This can explain why the coefficient estimate of in Column (1) of Table 5 is statistically insignificant, whereas it becomes statistically significant in Column (2).
-4.2. News Sharer Engagement: Lifetime and Post Count (RQ3 & RQ4)
-Summary statistics: Fig.
-3(d) shows that false news items have significantly longer lifetimes (mean of 1463.757 hours) than true news items (mean of 47.951 hours, , ).
+Given the second reason, we further examine the paired changes in topic diversity from the source news claims to the corresponding posts.
+In Fig.
+3(f), the positive coefficient estimate of is statistically significant ( , ).
+This suggests that the increase in topic diversity in false news from the claim level to the post level is significantly larger (0.659 standard deviations more) than in true news.
+This finding could partially explain the enlarged difference in topic diversity between true and false news posts, compared to their source claims.
+In addition, the coefficient estimate of in Fig.
+3(f) is significantly positive ( , ).
+This means that the topic diversity in false news that contains conspiracy theories increases 0.185 standard deviations more than that in false news without conspiracy theories from the claim level to the post level.
+This finding could partially explain why the coefficient estimate of in Fig.
+3(d) is not statistically significant, whereas it becomes statistically significant in Fig.
+3(e).
+In summary, we find that false news claims exhibit higher topic diversity compared to true news claims, and this difference is magnified at the post level (RQ1.1).
+Additionally, false news posts containing conspiracy theories have higher topic diversity than those without conspiracy theories, while no significant difference is observed at the claim level (RQ2.1).
+4.2. News Source Sharer Engagement: News Lifetime and Post Count (RQ1.2 & RQ2.2)
+4.2.1. Descriptive statistics.
+In terms of engagement at news source sharer level, we examine news lifetime that indicates the duration (in hours) from the creation time of the first post to the creation time of the last post for each news item on X.
+Fig.
+4(a) shows that false news items have significantly longer lifetimes (mean of 1463.757 hours) than true news items (mean of 47.951 hours, , ).
 CT news items have longer lifetimes (mean of 1968.141 hours) than false news items ( , ).
 Additionally, Fig.
-3(e) shows that false news claims receive more posts (mean of 38.015) than true news claims (mean of 3.159, , ).
-However, the number of posts received by CT news claims (mean of 44.976) has no statistically significant difference from that received by false news claims ( , ).
-Regression results: As shown in Column (1) of Table 6, the coefficient estimate of is significantly positive ( , ), which suggests that false news items have longer lifetimes to get new posts on X than true news items.
+4(b) shows that false news claims receive more posts (mean of 38.015) than true news claims (mean of 3.159, , ).
+However, the number of posts received by CT news items (mean of 44.976) does not differ significantly from that received by false news items ( , ).
+4.2.2. Regression results.
+We conduct regression analysis to further rigorously examine the effects of , , and on news lifetime and post count (Figs.
+4(c)–4(d)).
+In Fig.
+4(c), the coefficient estimate of is significantly positive ( , ).
+This suggests that, on average, false news items have lifetimes of 1.239 standard deviations longer for receiving posts on X compared to true news items.
 Additionally, the coefficient estimate of is also significantly positive ( , ).
-It means that false news items containing conspiracy theories have significantly longer lifetimes than false news items without conspiracy theories.
-However, the coefficient estimates for and are statistically insignificant.
-This suggests that topic diversity has no significant impact on the lifetime of news items, irrespective of the veracity.
-Additionally, in Column (2) of Table 6, the coefficient estimate of is significantly positive ( , ), indicating that the number of posts referencing a specific source news claim is significantly associated with its veracity.
-On average, false news claims are linked by times more posts on X compared to true news claims.
-Notably, the coefficient estimate of is not statistically significant for post count, which implies that the integration of conspiracy theories has no additional impact on the number of posts received by false news claims.
-The coefficient estimate of is statistically insignificant.
-This suggests that the topic diversity of a source news claim has no significant relationship with the number of online posts it receives.
-The reason for the enlarged difference in topic diversity between true and false news posts is not supported by this observation.
-Furthermore, the coefficient estimate for the interaction is also statistically insignificant, reinforcing the notion that topic diversity has no significant effect on post count, irrespective of the veracity of news claims.
-(1) | (2) | |
-Lifetime | Post Count | |
-Falsehood | ∗∗∗ | ∗∗∗ |
-Conspiracy | ∗ | |
-Diversity | ||
-Falsehood Diversity | ||
-Words | ∗∗∗ | ∗∗∗ |
-Emotions | ✓ | ✓ |
-Language | ✓ | ✓ |
-MonthYear | ✓ | ✓ |
-Intercept | ∗∗∗ | ∗∗∗ |
-0.481 | 0.188 |
-4.3. Post Viewer Engagement: Reposts, Likes, and Replies (RQ3 & RQ4)
-Summary statistics: Figs.
-3(f) – 3(h) reveal that, on average, true news posts get more engagement on X than false news posts in terms of reposts (true: 29.396, false: 14.403, , ), likes (true: 84.765, false: 45.234, , ), and replies (true: 6.294, false: 5.031, , ).
-However, it’s worth noting that the news posts that receive the most significant reposts, likes, or replies belong to the false category.
-Additionally, CT news posts receive more reposts (mean of 37.942, , ), likes (mean of 106.004, , ), and replies (mean of 25.173, , ) than overall false news posts.
-Notably, the difference in reply count between true and false news posts is the smallest compared to repost count and like count.
-Regression results: In Column (1) of Table 7, the coefficient estimate of is significantly negative ( , ), which means that false news posts receive fewer reposts than true news posts.
+This means that, on average, the lifetimes for false news items containing conspiracy theories are 0.452 standard deviations longer than that for false news items without conspiracy theories.
+However, the coefficient estimates for and are not statistically significant.
+This suggests that topic diversity has no significant impact on the lifetimes of news items, irrespective of their veracity.
+Additionally, in Fig.
+4(d), the coefficient estimate of is significantly positive ( , ), indicating that the number of posts referencing a specific source news item is significantly associated with its veracity.
+On average, false news items are linked by times more posts on X compared to true news items.
+Notably, the coefficient estimate of is not statistically significant for post count, which implies that the integration of conspiracy theories has no additional impact on the number of posts received by false news items.
+The coefficient estimate of is not statistically significant.
+It suggests that the topic diversity of a source news claim has no significant association with the number of online posts it receives.
+This finding does not support the first reason for the enlarged difference in topic diversity between true and false news posts, compared to that between true and false news claims (see Section 4.1.2).
+Furthermore, the coefficient estimate of the interaction term is also not statistically significant, reinforcing the finding that topic diversity has no significant effect on post count, irrespective of the veracity of news claims.
+In summary, we find that, at news source sharer level, misinformation with higher topic diversity does not receive more engagement in terms of news lifetime and post count (RQ1.2).
+However, false news items containing conspiracy theories tend to receive a higher number of posts on X compared to false news items without conspiracy theories but do not have significantly longer lifetimes (RQ2.2).
+4.3. Post Viewer Engagement: Reposts, Likes, and Replies (RQ1.2 & RQ2.2)
+4.3.1. Descriptive statistics.
+In terms of engagement at post viewer level, we examine the numbers of reposts, likes, and replies for news posts.
+To this end, we first conduct descriptive analysis without incorporating the defined control variables from content and poster characteristics.
+We find that, on average, true news posts receive more engagement on X than false news posts in terms of reposts (Fig.
+5(a); true: 29.396, false: 14.403, , ), likes (Fig.
+5(b); true: 84.765, false: 45.234, , ), and replies (Fig.
+5(c); true: 6.294, false: 5.031, , ).
+However, it is worth noting that news posts that receive the most significant reposts, likes, or replies belong to the false category.
+Additionally, CT news posts receive more reposts (mean of 37.942, , ) and replies (mean of 25.173, , ) than overall false news posts, while without statistically significant difference in like count ( , ).
+4.3.2. Regression results.
+Incorporating multiple control variables related to content and poster characteristics that can affect engagement with news posts, we conduct regression analysis to estimate the effects of , , and on the numbers of reposts, likes, and replies (Figs.
+5(d)–5(f)).
+In Fig.
+5(d), the coefficient estimate of is significantly negative ( , ), which means that false news posts receive fewer reposts than true news posts.
 On average, false news posts receive 22.6% fewer reposts than true news posts.
-Similarly, the coefficient estimate of in Column (2) is significantly negative ( , ).
-On average, false news posts receive 14.4% fewer likes than true news posts.
-However, the coefficient estimate of in Column (3) is significantly positive ( , ).
+Similarly, the coefficient estimate of in Fig.
+5(e) is significantly negative ( , ).
+This means that, on average, false news posts receive 14.4% fewer likes than true news posts.
+However, the coefficient estimate of in Fig.
+5(f) is significantly positive ( , ).
 This means that, on average, false news posts receive 49.2% more replies than true news posts.
-Together with the observations in Figs.
-3(f) – 3(h), the results indicate that the interaction of reply with the veracity of posts may vary from the interactions of reposting and liking.
-Additionally, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive.
-It indicates that heightened topic diversity is associated with increased online engagement in terms of reposts, likes, and replies.
-However, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly negative.
-This suggests that the positive association of topic diversity with engagement on X is reduced within the context of false news posts.
-On the contrary, the coefficient estimates of in Columns (1) ( , ), (2) ( , ), and (3) ( , ) are significantly positive.
-This means that the integration of conspiracy theories is positively linked to the engagement with false news posts on X at post viewer level.
+Additionally, the coefficient estimates of in Fig.
+5(d) ( , ), Fig.
+5(e) ( , ), and Fig.
+5(f) ( , ) are significantly positive.
+This indicates that heightened topic diversity is associated with increased online engagement in reposts, likes, and replies.
+Specifically, a one standard deviation increase in topic diversity is associated with 19.8% more reposts, 21.2% more likes, and 14.7% more replies.
+However, the coefficient estimates of in Fig.
+5(d) ( , ), Fig.
+5(e) ( , ), and Fig.
+5(f) ( , ) are significantly negative.666The coefficient estimates of across repost count, like count, and reply count remain significantly positive in the subset analysis for false news posts.
+See Table S4 in Suppl.
+S1 for full estimation results.
+This suggests that the positive association of topic diversity with engagement on X is not significantly stronger within the context of false news posts compared to true news posts.
+On the contrary, the coefficient estimates of in Fig.
+5(d) ( , ), Fig.
+5(e) ( , ), and Fig.
+5(f) ( , ) are significantly positive.
+This means that the integration of conspiracy theories within false news posts is positively linked to engagement at post viewer level.
 Specifically, false news posts that contain conspiracy theories receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
-(1) | (2) | (3) | |
-Repost Count | Like Count | Reply Count | |
-Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
-Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
-Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
-Emotions | ✓ | ✓ | ✓ |
-Language | ✓ | ✓ | ✓ |
-MonthYear | ✓ | ✓ | ✓ |
-Intercept | ∗∗∗ | ∗∗∗ | |
-News-level REs | ✓ | ✓ | ✓ |
+Clarification for the discrepancies between descriptive statistics and regression results: We observe two discrepancies between descriptive analysis and regression analysis: (i) In the descriptive analysis, Fig.
+5(c) shows that true news posts receive more replies than false news posts, whereas Fig.
+5(f) indicates that false news posts receive more replies than true news posts.
+(ii) No significant difference between CT news posts and overall false news posts is observed in the CCDF of like count (Fig.
+5(b)), while Fig.
+5(e) indicates that false news posts containing conspiracy theories receive more likes than those without conspiracy theories.
+These discrepancies suggest that the falsehood may correlate with other variations in the control variables, potentially complicating the findings in the descriptive analysis and underscoring the necessity of incorporating control variables that can affect engagement.
+For instance, emotions, included as control variables in our regression models, have been well documented as a significant driver in facilitating online post sharing (Vosoughi et al., 2018; Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b; Solovev and Pröllochs, 2022).
+We find that falsehood is positively correlated with anger ( , ) and surprise ( , ) but negatively correlated with joy ( , ; see Table S6 in Suppl.
+S2).
+This suggests that false news posts express more anger and surprise but less joy compared to true news posts.
+Within the estimation results of our regression models, the coefficient estimates of for repost count ( , ), like count ( , ), and reply count ( , ) are consistently positive and statistically significant (see Table S3 in Suppl.
+S1 for full estimation results).
+Similarly, the coefficient estimates of for repost count ( , ), like count ( , ), and reply count ( , ) are also consistently positive and statistically significant.
+Our results suggest that news posts expressing higher levels of anger or surprise tend to receive more reposts, likes, and replies, aligning with findings from previous research (Vosoughi et al., 2018; Chuai and Zhao, 2022).
+However, the coefficient estimate of is not statistically significant for repost count, while significantly positive for reply count ( , ).
+This suggests that news posts expressing more joy are more likely to receive replies but do not receive more reposts.
+In light of this, we find two opposite effects: falsehood contributes to an increase in replies for false news posts, while the lower expression of joy in false news posts, compared to true news posts, links to fewer replies.
+The aggregation of these two opposite effects in the descriptive analysis could partially explain the discrepancy observed between Fig.
+5(c) and Fig.
+5(f).
+In summary, at post viewer level, news posts with higher topic diversity receive more reposts, likes, and replies; however, this association is not stronger for false news posts (RQ 1.2).
+Additionally, false news posts containing conspiracy theories receive more reposts, likes, and replies compared to false news posts without conspiracy theories (RQ 2.2).
 4.4. Robustness Checks
-To ensure the robustness of our findings, we conduct several additional checks.
-(i) The cross-correlations among the independent variables are reported in Table S1 (news level) and Table S2 (post level).
-The variance inflation factors (VIFs) are also reported in Table S3.
-All the VIFs are significantly lower than the threshold of 4, which indicates that multicollinearity is not an issue for our analysis.
-(ii) We recalculate topic diversity at the theme level and repeat our analysis (Table S4 – Table S6).
-(iii) We further balance the samples between true and false news items through propensity score matching and reduce potential unobserved confounding factors (Table S7 – Table S9).
-(iiii) We use an alternative topic modeling method (BERTopic) to cluster topics again and repeat our analysis (Table S10 – Table S12).
+To ensure the robustness of our findings, we conduct several additional checks: (i) We examine the variance inflation factors (VIFs) among the independent variables at both news claim and news post levels (see Table S7 in Suppl.
+S2).
+All the VIFs are significantly lower than the critical threshold of 4, which indicates that multicollinearity is not an issue for our analysis.
+(ii) To test whether our findings are robust with potential noise in topic keywords, we conduct our analysis again using the original 26 topics and associated keywords without manual validation (see details in Suppl.
+S3).
+(iii) To ensure that our findings are not influenced by the topic granularity, we recalculate topic diversity at the theme level and repeat our analysis (see details in Suppl.
+S4).
+(iv) We further balance the samples between true and false news items through propensity score matching and reduce potential unobserved confounding factors (see details in Suppl.
+S5).
+(v) We extend conspiracy theories by including QAnon and repeat our analysis (see details in Suppl.
+S6).777QAnon is a far-right conspiracy theory alleging that a cabal of politicians and media personalities are engaged in a war to undermine society (Papasavva et al., 2022; Bär et al., 2023; Kim and Kim, 2023; Moore et al., 2024).
+The QAnon discourse has become more diverse and evolved into a global phenomenon during the COVID-19 pandemic (Hoseini et al., 2023).
+We adopt a list of QAnon-specific keywords from previous research (Sharma et al., 2022a) to identify QAnon conspiracy theory in our dataset.
+The performance of this keyword list in identifying QAnon discourse has been manually validated, showing high agreement with human assessments (Bär et al., 2023).
+As a result, we identify 6 QAnon news claims and 302 QAnon news posts.
+(vi) Given that news posts can also debunk linked false news items, we exclude potential debunking posts and repeat our analysis (see details in Suppl.
+S7).
+(vii) We use an alternative topic modeling method (BERTopic) to cluster topics again and repeat our analysis (see details in Suppl.
+S8).
 All the results are robust and consistently support our main findings.
-5. Discussion
-The spread of misinformation surrounding health crises is a serious public concern.
-In this paper, we present a comprehensive analysis of the impact of topic diversity and conspiracy theories on the engagement dynamics of COVID-19 misinformation.
-Specifically, we examine user engagement at two levels spanning from the origin as source news items to their propagation through online posts and subsequent social engagement (i.
-e., reposts, likes, and replies) on X.
+4.5. Summary of Main Findings
 Our analysis yields several key findings: (i) False news, especially when accompanied by conspiracy theories, exhibits higher topic diversity compared to true news.
 This discrepancy increases from source news claims to news posts on X.
-(ii) At news sharer level, neither the news lifetime nor the number of news posts is significantly associated with topic diversity.
+(ii) At news source sharer level, neither the news lifetime nor the number of news posts is significantly associated with topic diversity.
 However, false news has a longer lifetime and receives more posts compared to true news.
 Additionally, the integration of conspiracy theories is associated with a longer lifetime for false news but has no significant association with the number of false news posts.
-(iii) At post viewer level, posts with higher topic diversity receive more reposts, likes, and replies.
-However, the positive effect of topic diversity is moderated by the news veracity, with higher topic diversity being linked to more engagement with true news posts compared to false news posts.
-(iiii) The integration of conspiracy theories is linked to more engagement with COVID-19 misinformation at post viewer level.
+(iii) At post viewer level, posts with higher topic diversity receive more reposts, likes, and replies, while this association is not stronger in false news posts.
+(iv) The integration of conspiracy theories is linked to more engagement with COVID-19 misinformation at post viewer level.
 False news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories.
-These findings have useful implications for both research and practical applications.
-Implications: First, the observed high topic diversity in COVID-19 misinformation indicates that the interconnection of multiple topics becomes a characteristic of misinformation narratives during health crises.
-Previous research also reveals a coordinated effort among misinformation websites from multiple domains to hyperlink to each other (Hanley et al., 2023).
-Notably, the topic diversity in COVID-19 misinformation significantly increases from its source claims to online posts.
-This mutation may contribute to the adaptability of misinformation, allowing it to resonate with diverse audiences and contexts (Yan et al., 2022).
-Additionally, the finding that misinformation containing conspiracy theories exhibits the strongest topic diversity aligns with previous research (Miani et al., 2022), highlighting the interconnected nature of conspiracy narratives.
-Within the context of multi-topic misinformation, individuals tend to adaptively connect related information and make associative inferences to reduce uncertainty and gain control in a way that is congruent to their prior knowledge (Lee et al., 2023; Miani et al., 2022).
+5. Discussion
+5.1. Relevance
+The spread of misinformation surrounding health crises is a serious public concern, intersecting multiple domains in society.
+Here, we present an empirical analysis of fact-checked source news claims and their associated posts on X to explore how topic diversity and conspiracy theories influence engagement with multi-topic misinformation during the COVID-19 pandemic.
+Understanding the characteristics of misinformation and their effects on user engagement is a crucial and growing area of research for developing effective countermeasures and mitigating the spread of misinformation (Vosoughi et al., 2018; Chuai and Zhao, 2022; Chuai et al., 2024b; Ecker et al., 2022, 2024).
+Our study makes two main contributions: (i) We analyze the characteristics of misinformation and its spread from the perspectives of topic diversity and conspiracy theories; (ii) We examine how these characteristics shape engagement with misinformation at two distinct user levels, i.
+e., news source sharer level and post viewer level.
+Specifically, we investigate online user engagement from the shares of source news items by news source sharers to the subsequent engagement with news posts by post viewers on X (i.
+e., reposts, likes, and replies).
+Our findings offer valuable insights for further research and practical strategies for combating the spread of online misinformation during health crises and beyond.
+5.2. Research Implications
+5.2.1. Heightened topic diversity in misinformation/conspiracy theories.
+The observed high topic diversity in COVID-19 misinformation suggests that the interconnection of multiple topics becomes a characteristic of misinformation narratives during health crises.
+Here, the topic diversity reflects the interconnectedness of various topics within a single piece of misinformation, potentially enhancing its novelty.
+Different from our study, Vosoughi et al.
+(2018) investigated the topic similarity between users’ reposted true (or false) posts and their previously viewed posts.
+They found that false posts are more dissimilar from previously viewed posts compared to true posts.
+However, given users’ topic interests and recommendation algorithms’ tendency to promote similar content, users potentially consume more content that aligns with their interests (Cinus et al., 2022; Engel et al., 2023).
+These seemingly conflicting findings could be partially reconciled by our insights on topic diversity: misinformation may strategically link familiar topics together with diverse, novel topics, thereby drawing users’ attention to new information and increasing their susceptibility.
+Particularly, our finding that misinformation containing conspiracy theories exhibits the strongest topic diversity underscores this deliberate strategy in conspiracy narratives.
+Moreover, previous research reveals coordinated efforts among misinformation websites from multiple domains to hyperlink to each other, further supporting our interpretation (Hanley et al., 2023).
+Additionally, we find that the topic diversity in COVID-19 misinformation significantly increases from its source claims to subsequent online posts.
+This mutation/evolution may contribute to the adaptability of misinformation, allowing it to resonate with diverse audiences and contexts (Yan et al., 2022).
+For instance, QAnon is a far-right political conspiracy theory that originated in 2017.
+It became more diverse during COVID-19, evolving and expanding beyond political themes to incorporate pandemic-related topics (Hoseini et al., 2023).
+Theoretically, within the context of multi-topic misinformation, individuals tend to adaptively connect related information and make associative inferences to reduce uncertainty and gain control in a way that is congruent to their prior knowledge (Lee et al., 2023; Miani et al., 2022).
 Understanding this cognitive process is essential for developing targeted interventions aimed at reducing susceptibility to multi-topic misinformation, especially when accompanied by conspiracy theories.
-Second, at the news sharer level, the observed longer lifetime and higher post count on X for false news items, compared to true news items, indicate the endurance and prevalence of misinformation on social media during COVID-19.
-Notably, few prior studies focus on engagement at news sharer level and study how false news enters the social media platforms via news sharers.
-The extended lifetime we observe in this paper may provide false news with more opportunities to accumulate engagement on social media, potentially reaching a broader audience.
-Moreover, the lifetime and post count appear to be largely attributed to the veracity of the news claims, as opposed to other variables including and .
-Given this, it is plausible that the prolonged lifetime and high post count for false news claims mainly result from deliberately planned dissemination strategies by misinformation providers, involving means such as social bots and monetization (Teng et al., 2022; Moffitt et al., 2021; Ballard et al., 2022).
-Consequently, the role of individuals’ self-motivated sharing behaviors in the endurance and prevalence of COVID-19 misinformation is limited at news sharer level.
-However, massive posts repeating the same false claim can still increase individuals’ susceptibility and beliefs.
-In this regard, social media platforms should take measures to restrict repeated original posts from the same false claims and their mutations, which is a lack of discussion in the previous research.
-Third, as opposed to the engagement at news sharer level, the engagement at post viewer level (i.
-e., reposts, likes, and replies) on X might be largely driven by human users rather than social bots during COVID-19 (Teng et al., 2022).
-Organic user engagement from human users can be influenced by multiple factors.
-For instance, our results in Table 7 show that the engagement at the post viewer level can be affected by factors related to content (e.
-g., topic diversity, conspiracy theories, emotions) and context (e.
-g., verified status and follower count).
-Previous findings on the virality of misinformation at post viewer level are mixed (Vosoughi et al., 2018; Drolsbach and Pröllochs, 2023).
-We also find that false news posts receive fewer reposts and likes but receive more replies, compared to true news posts.
-Of note, the veracity of information is not explicit to users before being verified.
-Therefore, there might be other content or context factors that are related to the veracity and foster engagement.
-While a large body of research has examined the engagement at the posts viewer level (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b, a; Wang et al., 2023a; Solovev and Pröllochs, 2022), the understanding on the role of topic diversity and conspiracy theories in shaping engagement of post viewers with misinformation is still missing.
-Our findings show that higher topic diversity is associated with more reposts, likes, and reposts for both true and false COVID-19 news posts on X.
-However, despite COVID-19 misinformation having higher topic diversity, the positive effect of topic diversity on its engagement is reduced, requiring further detailed study.
-Notably, conspiracy theories can significantly enhance engagement with COVID-19 misinformation.
-To efficiently reduce social engagement with (health) misinformation, social media platforms should exclusively focus on the typical misinformation that exhibits high topic diversity and contains conspiracy theories.
-Limitations: Our empirical study reports associations between variables and avoids making causal claims.
-Future research could benefit from experimental designs to establish causal relationships and better understand the influence of topic diversity and conspiracy theories on engagement with online misinformation during health crises.
-Moreover, our analysis relies on data from X, limiting the generalizability of our findings to other social media platforms.
-Though the prevalence of health misinformation was the highest on X during COVID-19 (Suarez-Lledo and Alvarez-Galvez, 2021), different platforms may exhibit unique engagement patterns and user behaviors, necessitating further research across the platforms to capture a comprehensive understanding of online misinformation dynamics during health crises.
-Additionally, the absence of detailed engagement timelines at post viewer level in our dataset hinders a more granular analysis of how topic diversity and conspiracy theories influence engagement with COVID-19 misinformation over time on social media.
+Taken together, our findings underscore the importance of considering the multi-topic nature of misinformation in further research.
+5.2.2. Engagement patterns at news source sharer level and post viewer level.
+During health crises marked by high uncertainty, novel information can update users’ understanding and knowledge of the current state, helping them manage anxiety and fostering their engagement intentions (Vosoughi et al., 2018; Freiling et al., 2023; Chuai and Zhao, 2022; Chen et al., 2023).
+Furthermore, sharing novel posts can make users feel that they possess rare and important information that others may lack, thereby boosting their self-esteem (Douglas et al., 2019; Chuai and Zhao, 2022).
+In this light, our findings suggest that the high topic diversity and the incorporation of conspiracy theories in misinformation can enhance its perceived novelty, thereby driving engagement.
+However, through further examining how topic diversity and conspiracy theories shape engagement at news source sharer level and post viewer level, we find distinct engagement patterns at the two user levels.
+Engagement patterns at news source sharer level: At news source sharer level, the observed longer lifetime and higher post count on X for false news items, compared to true news items, indicate the persistence and widespread reach of misinformation on social media during COVID-19.
+The persistence of false news items provides them with more opportunities to accumulate engagement on social media, potentially reaching a broader audience.
+Moreover, the lifetime and post count appear to be largely attributed to the veracity of news claims, as opposed to other variables including and .
+Given this, it is plausible that the prolonged lifetime and high post count for false news items mainly result from deliberately planned dissemination strategies by misinformation providers, involving means such as social bots and monetization (Teng et al., 2022; Moffitt et al., 2021; Ballard et al., 2022).
+This perspective also aligns with the prior finding that social bots amplify low-credibility content during the initial spreading stages (Shao et al., 2018).
+Consequently, the role of individuals’ self-motivated sharing behaviors in the persistence and diffusion of COVID-19 misinformation is limited at news source sharer level.
+However, massive posts repeating the same false claims with adaptability – i.
+e., through increased topic diversity – can still heighten individuals’ susceptibility and beliefs.
+In this regard, social media platforms should take measures to restrict repeated posts linking to the same false news items and their mutations.
+Engagement patterns at post viewer level: As opposed to engagement at news source sharer level, engagement at post viewer level (i.
+e., reposts, likes, and replies) on X might be largely driven by organic interactions from human users rather than social bots during COVID-19 (Teng et al., 2022).
+Accordingly, we find that topic diversity and conspiracy theories become significant in driving engagement with misinformation at post viewer level in terms of reposts, likes, and replies.
+To efficiently reduce human users’ engagement with misinformation at this level, social media platforms need to prioritize targeting misinformation that exhibits high topic diversity and incorporates conspiracy theories.
+Additionally, our analysis reveals that replies differ from reposts and likes in their engagement with misinformation.
+Specifically, excluding the effects of topic diversity, conspiracy theories, and other control variables, false news posts receive more replies but fewer reposts and likes than true news posts.
+Distinct from reposts and likes, replies can not only support but also debunk false news posts.
+Given this, the higher number of replies for false news posts than true news posts can be triggered by debunking misinformation (Zannettou, 2021).
+Our supplementary analysis of quotes shows that, although not statistically significant, false news posts also have a trend to receive more quotes than true news posts (see Suppl.
+S9 for full estimation results).
+This finding suggests that quotes serve a similar function as replies.
+Taken together, our findings suggest that engagement with misinformation might be primarily driven by social bots at news source sharer level, while human users dominate engagement at post viewer level.
+This distinction highlights the need to examine misinformation engagement and develop intervention strategies at a multi-level framework.
+5.3. Practical Implications
+5.3.1. Implementing cross-platform fact-checking.
+Currently, social media providers are actively implementing measures to control the spread of misinformation on their platforms, including third-party professional fact-checking by experts and community-based fact-checking driven by crowds (Chuai et al., 2024b; Ecker et al., 2022).
+However, these content moderation policies/features primarily function within individual platforms.
+Our findings suggest that misinformation from external sources, even if fact-checked on one platform, can still enter another platform through a large volume of posts.
+To prevent this spillover of misinformation across platforms, coordinated efforts among social media platforms, such as sharing fact-checking information, are needed.
+5.3.2. Expanding fact-checks to misinformation with same links.
+Given that misinformation can be repeatedly shared on social media platforms through the same source links, an effective strategy is to directly moderate posts that contain previous fact-checked misinformation links.
+For example, X has rolled out a Community Notes feature that allows users to write fact-checking notes as additional context displayed on the corresponding misleading posts.888More details are available at https://communitynotes.x.com/guide/en/about/introduction This community-based fact-checking approach has proven effective in reducing the spread of misleading posts (Chuai et al., 2024a).
+A recent update to Community Notes on X ensures that helpful notes that indicate posts’ media elements, such as images or videos, are misleading will be visible on all posts containing the same media.
+Similarly, fact-checking notes about misinformation links could be expanded to cover all posts that contain the same links.
+5.3.3. Empowering logical reasoning and critical thinking for social media users.
+To reduce users’ susceptibility to misinformation with high topic diversity and promote rational information processing, tools that can enhance users’ logical reasoning and critical thinking abilities should be adopted.
+The rise of Large Language Models (LLMs) presents a promising opportunity to help users conduct logical reasoning and identify potentially conflicting or fictional information.
+Although current LLMs may still have limitations in reasoning tasks, particularly for misinformation identification (Pareek et al., 2024), this approach represents a promising direction as LLMs continue to improve (Zavolokina et al., 2024).
+5.4. Limitations and Further Research
+Our study is subject to several limitations.
+First, our analysis is based on real-world social media data and, as is typical with observational studies, is limited to reporting associations between variables rather than causal claims.
+Future research could benefit from experimental designs and user studies to establish causal relationships and explore heterogeneity in the effects of topic diversity and conspiracy theories on engagement across different user groups.
+Second, our findings on engagement can be influenced by social media recommendation algorithms and potential soft moderation implemented on misleading posts.
+Specifically, while our study mainly focuses on user-driven engagement at two levels and discusses the potential role of social bots, social media algorithms can selectively recommend posts and boost their visibility, potentially influencing user engagement and adding complexity to our results.
+Given that posts linking to the same news items can be treated similarly by recommendation algorithms or have similar audiences, the incorporation of news-specific random effects, i.
+e., random intercepts, in our models helps mitigate this concern.
+However, the precise role of recommendation algorithms warrants further exploration in future research.
+Additionally, due to the absence of information indicating whether and when soft moderation was applied to certain misleading posts, we cannot control for the potential effect of soft moderation on subsequent engagement.
+Previous research has shown that soft moderation interventions – e.
+g., attaching warning tags – adopted by X during COVID-19 can reduce belief in misinformation and thus lower the spread/visibility of misleading posts (Sharevski et al., 2022).
+Moreover, a recent study demonstrates that community-based fact-checking approach is effective in reducing the spread of misleading posts on X (Chuai et al., 2024a).
+With the potential impact of such interventions in reducing engagement, we still find that false news has a longer lifetime than true news.
+This indicates that our results are not overestimated.
+On the other hand, prior research suggests that warning labels can also prompt interactions aimed at debunking false claims (Zannettou, 2021).
+To address this concern, we exclude potential debunking posts based on debunk-specific keywords and conduct our analysis again.
+Our results are robust and consistently support our main findings.
+Furthermore, the absence of detailed engagement timelines at post viewer level in our dataset hinders a more granular analysis of how topic diversity and conspiracy theories influence engagement with misinformation over time on social media.
+This limitation highlights an important area for further examination.
+Finally, our analysis is based on a single curated public dataset that includes social context on one platform – X (Li et al., 2020).
+Although this dataset offers unique features, such as multilingual content, it may introduce potential selection bias during data collection.
+To the best of our knowledge, no other publicly available datasets are suitable for addressing our research questions.
+Moreover, current restrictions on social media platform APIs make it increasingly challenging for researchers to access comprehensive data.
+We advocate for greater openness from social media platforms to facilitate academic research and encourage further studies to validate and generalize our findings across broader datasets and other social media platforms.
 6. Conclusion
@@
 It is crucial to understand the factors and mechanisms that contribute to the engagement with health misinformation on social media.
-Here, we provide a comprehensive analysis of how topic diversity and conspiracy theories shape engagement with COVID-19 misinformation from news sharer level to post viewer level.
+Here, we provide a comprehensive analysis of how topic diversity and conspiracy theories shape engagement with COVID-19 misinformation from news source sharer level to post viewer level.
 We find that high topic diversity and the integration of conspiracy theories are characteristics of COVID-19 misinformation.
-Topic diversity has no significant association with the engagement at news sharer level.
+Topic diversity has no significant association with the engagement at news source sharer level.
 However, topic diversity and conspiracy theories can significantly enhance engagement with COVID-19 misinformation at post viewer level.
-Our findings provide insights into understanding the engagement with health misinformation and highlight the engagement patterns for news sharers and post viewers regarding topic diversity and conspiracy theories.
+Our findings provide insights into understanding the engagement with health misinformation and highlight the engagement patterns for news source sharers and post viewers regarding topic diversity and conspiracy theories.
 These insights are valuable for developing targeted interventions at both user levels during health crises.
@@
 Conspiracy brokers: Understanding the monetization of YouTube conspiracy theories.
-In WWW.
+In Proceedings of the ACM Web Conference 2022.
+2707–2718.
+- Bär et al.
+(2023) Dominik Bär, Nicolas Pröllochs, and Stefan Feuerriegel.
+2023.
+Finding Qs: Profiling QAnon supporters on Parler.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+17.
+34–46.
 - Chandrasekaran et al.
@@
 ArXiv.
-- Chen et al. (2023) Lucia L. Chen, Steven R. Wilson, Sophie Lohmann, and Daniela V. Negraia. 2023. What are you anxious about? Examining subjects of anxiety during the COVID-19 pandemic. In ICWSM.
-- Chuai et al. (2022) Yuwei Chuai, Yutian Chang, and Jichang Zhao. 2022. What really drives the spread of COVID-19 tweets: A revisit from perspective of content. In DSAA.
-- Chuai et al. (2024) Yuwei Chuai, Haoye Tian, Nicolas Pröllochs, and Gabriele Lenzini. 2024. Did the roll-out of Community Notes reduce engagement with misinformation on X/Twitter?. In CSCW.
+- Chen et al.
+(2023) Lucia L.
+Chen, Steven R.
+Wilson, Sophie Lohmann, and Daniela V.
+Negraia.
+2023.
+What are you anxious about?
+Examining subjects of anxiety during the COVID-19 pandemic.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+17.
+137–148.
+- Chuai et al.
+(2022) Yuwei Chuai, Yutian Chang, and Jichang Zhao.
+2022.
+What really drives the spread of COVID-19 tweets: A revisit from perspective of content.
+In 2022 IEEE 9th International Conference on Data Science and Advanced Analytics.
+1–10.
+- Chuai et al.
+(2024a) Yuwei Chuai, Moritz Pilarski, Thomas Renault, David Restrepo-Amariles, Aurore Troussel-Clément, Gabriele Lenzini, and Nicolas Pröllochs.
+2024a.
+Community-based fact-checking reduces the spread of misleading posts on social media.
+ArXiv.
+- Chuai et al.
+(2024b) Yuwei Chuai, Haoye Tian, Nicolas Pröllochs, and Gabriele Lenzini.
+2024b.
+Did the roll-out of Community Notes reduce engagement with misinformation on X/Twitter?
+Proceedings of the ACM on Human-Computer Interaction 8, CSCW2 (2024), 1–52.
 - Chuai and Zhao (2022) Yuwei Chuai and Jichang Zhao. 2022. Anger can make fake news viral online. Frontiers in Physics 10 (2022), 970174.
-- Chuai et al. (2023) Yuwei Chuai, Jichang Zhao, Nicolas Pröllochs, and Gabriele Lenzini. 2023. Political Elites in False Statements on the Internet. ArXiv.
+- Chuai et al.
+(2025) Yuwei Chuai, Jichang Zhao, Nicolas Pröllochs, and Gabriele Lenzini.
+2025.
+Is fact-checking politically neutral?
+Asymmetries in how U.S.
+fact-checking organizations pick up false statements mentioning political elites.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+19.
+403–429.
+- Cinelli et al.
+(2021) Matteo Cinelli, Gianmarco De Francisci Morales, Alessandro Galeazzi, Walter Quattrociocchi, and Michele Starnini.
+2021.
+The echo chamber effect on social media.
+Proceedings of the National Academy of Sciences 118, 9 (2021), e2023301118.
+- Cinus et al.
+(2022) Federico Cinus, Marco Minici, Corrado Monti, and Francesco Bonchi.
+2022.
+The effect of people recommenders on echo chambers and polarization.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+16.
+90–101.
 - Davies et al.
@@
 - Douglas and Sutton (2018) Karen M. Douglas and Robbie M. Sutton. 2018. Why conspiracy theories matter: A social psychological analysis. European Review of Social Psychology 29, 1 (2018), 256–298.
-- Drolsbach and Pröllochs (2023) Chiara Drolsbach and Nicolas Pröllochs. 2023. Diffusion of community fact-checked misinformation on Twitter. In CSCW.
+- Douglas et al.
+(2019) Karen M Douglas, Joseph E Uscinski, Robbie M Sutton, Aleksandra Cichocka, Turkay Nefes, Chee Siang Ang, and Farzin Deravi.
+2019.
+Understanding conspiracy theories.
+Political Psychology 40 (2019), 3–35.
+- Drolsbach and Pröllochs (2023) Chiara Drolsbach and Nicolas Pröllochs.
+2023.
+Diffusion of community fact-checked misinformation on Twitter.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (2023), 1–22.
+- Ecker et al.
+(2024) Ullrich Ecker, Jon Roozenbeek, Sander van der Linden, Li Qian Tay, John Cook, Naomi Oreskes, and Stephan Lewandowsky.
+2024.
+Misinformation poses a bigger threat to democracy than you might think.
+Nature 630, 8015 (2024), 29–32.
 - Ecker et al.
@@
 PLoS ONE 17, 10 (2022), e0276082.
+- Engel et al.
+(2023) Kristen Engel, Shruti Phadke, and Tanushree Mitra.
+2023.
+Learning from the ex-believers: Individuals’ journeys in and out of conspiracy theories online.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (2023), 1–37.
 - Erokhin et al. (2022) Dmitry Erokhin, Abraham Yosipof, and Nadejda Komendantova. 2022. COVID-19 conspiracy theories discussion on Twitter. Social Media + Society 8, 4 (2022), 20563051221126051.
@@
 Growing polarization around climate change on social media.
-Nature Climate Change (2022), 1–8.
+Nature Climate Change 12, 12 (2022), 1114–1121.
 - Farrell (2019) Justin Farrell. 2019. The growth of climate change misinformation in US philanthropy: evidence from natural language processing. Environmental Research Letters 14, 3 (2019), 034013.
@@
 (2023) Isabelle Freiling, Nicole M Krause, Dietram A Scheufele, and Dominique Brossard.
-2021.
+2023.
 Believing and sharing misinformation, fact-checks, and accurate information on social media: The role of anxiety during COVID-19.
-New Media & Society (2021), 14614448211011451.
+New Media & Society 25, 1 (2023), 141–162.
 - Garrett and Bond (2021) R. Kelly Garrett and Robert M. Bond. 2021. Conservatives’ susceptibility to political misperceptions. Science Advances 7, 23 (2021), eabf1234.
@@
 A golden age: Conspiracy theories’ relationship with misinformation outlets, news media, and the wider Internet.
-In CSCW.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (2023), 1–33.
 - Hart et al. (2020) P. Sol Hart, Sedona Chinn, and Stuart Soroka. 2020. Politicization and polarization in COVID-19 news coverage. Science Communication 42, 5 (2020), 679–697.
@@
 Journal of Management Information Systems 38, 4 (2021), 1039–1066.
+- Hoseini et al.
+(2023) Mohamad Hoseini, Philipe Melo, Fabricio Benevenuto, Anja Feldmann, and Savvas Zannettou.
+2023.
+On the globalization of the QAnon conspiracy theory through Telegram.
+In Proceedings of the 15th ACM Web Science Conference 2023.
+75–85.
 - Huang et al.
@@
 - Jenke (2023) Libby Jenke. 2023. Affective polarization and misinformation belief. Political Behavior (2023), 1–60.
-- Kim and Kim (2023) Soojong Kim and Jisu Kim. 2023. The information ecosystem of conspiracy theory: Examining the QAnon narrative on Facebook. (2023).
-- Kou et al. (2017) Yubo Kou, Xinning Gui, Yunan Chen, and Kathleen Pine. 2017. Conspiracy talk on social media: Collective sensemaking during a public health crisis. In CSCW.
+- Kim and Kim (2023) Soojong Kim and Jisu Kim.
+2023.
+The information ecosystem of conspiracy theory: Examining the QAnon narrative on Facebook.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (2023), 1–24.
+- Kou et al.
+(2017) Yubo Kou, Xinning Gui, Yunan Chen, and Kathleen Pine.
+2017.
+Conspiracy talk on social media: Collective sensemaking during a public health crisis.
+Proceedings of the ACM on Human-Computer Interaction 1, CSCW (2017), 1–21.
 - Kreps and Kriner (2020) S.
@@
 Science Advances 6, 43 (2020), eabd4563.
-- Lee et al. (2023) Sian Lee, Haeseung Seo, Dongwon Lee, and Aiping Xiong. 2023. Associative inference can increase people’s susceptibility to misinformation. In ICWSM.
+- Lee et al.
+(2023) Sian Lee, Haeseung Seo, Dongwon Lee, and Aiping Xiong.
+2023.
+Associative inference can increase people’s susceptibility to misinformation.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+17.
+530–541.
 - Lewandowsky (2021) Stephan Lewandowsky. 2021. Climate change disinformation and how to combat it. Annual Review of Public Health 42 (2021), 1–21.
@@
 - Moffitt et al. (2021) J. D. Moffitt, Catherine King, and Kathleen M. Carley. 2021. Hunting conspiracy theories during the COVID-19 pandemic. Social Media + Society 7, 3 (2021), 20563051211043212.
+- Moore et al.
+(2024) Ryan C Moore, Ross Dahlke, Peter L Forberg, and Jeffrey T Hancock.
+2024.
+The private life of QAnon: A mixed methods investigation of Americans’ exposure to QAnon content on the web.
+Proceedings of the ACM on Human-Computer Interaction 8, CSCW2 (2024), 1–34.
 - Ngai et al.
@@
 The impact of Twitter labels on misinformation spread and user engagement: Lessons from Trump’s election tweets.
-In WWW.
+In Proceedings of the ACM Web Conference 2022.
+2541–2551.
+- Papasavva et al.
+(2022) Antonis Papasavva, Max Aliapoulios, Cameron Ballard, Emiliano De Cristofaro, Gianluca Stringhini, Savvas Zannettou, and Jeremy Blackburn.
+2022.
+The Gospel according to Q: Understanding the QAnon conspiracy from the perspective of canonical information.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+16.
+735–746.
+- Pareek et al.
+(2024) Saumya Pareek, Niels van Berkel, Eduardo Velloso, and Jorge Goncalves.
+2024.
+Effect of explanation conceptualisations on trust in AI-assisted credibility assessment.
+Proceedings of the ACM on Human-Computer Interaction 8, CSCW2 (2024), 1–31.
 - Pierri et al.
@@
 presidential election.
-In CSCW.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (2023), 1–39.
 - Pröllochs et al.
@@
 Hyperpartisanship, disinformation and political conversations on Twitter: The Brazilian presidential election of 2018.
-In CSCW.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+14.
+569–578.
 - Robertson et al.
@@
 Royal Society open science 7, 10 (2020), 201199.
-- Schmitz et al. (2023) Matheus Schmitz, Goran Muric, and Keith Burghardt. 2023. Detecting anti-vaccine users on Twitter. In ICWSM.
-- Sharma et al. (2022) Karishma Sharma, Yizhou Zhang, and Yan Liu. 2022. COVID-19 vaccine misinformation campaigns and social media narratives. In ICWSM.
-- Solovev and Pröllochs (2022) Kirill Solovev and Nicolas Pröllochs. 2022. Moral emotions shape the virality of COVID-19 misinformation on social media. In WWW.
+- Schmitz et al.
+(2023) Matheus Schmitz, Goran Muric, and Keith Burghardt.
+2023.
+Detecting anti-vaccine users on Twitter.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+17.
+787–795.
+- Sebo and de Lucia (2024) Paul Sebo and Sylvain de Lucia.
+2024.
+Performance of machine translators in translating French medical research abstracts to English: A comparative study of DeepL, Google Translate, and CUBBITT.
+PLoS ONE 19, 2 (2024), e0297183.
+- Shao et al.
+(2018) Chengcheng Shao, Giovanni Luca Ciampaglia, Onur Varol, Kai-Cheng Yang, Alessandro Flammini, and Filippo Menczer.
+2018.
+The spread of low-credibility content by social bots.
+Nature Communications 9, 1 (2018), 1–9.
+- Sharevski et al.
+(2022) Filipo Sharevski, Raniem Alsaadi, Peter Jachim, and Emma Pieroni.
+2022.
+Misinformation warnings: Twitter’s soft moderation effects on COVID-19 vaccine belief echoes.
+Computers & Security 114 (2022), 102577.
+- Sharma et al.
+(2022a) Karishma Sharma, Emilio Ferrara, and Yan Liu.
+2022a.
+Characterizing online engagement with disinformation and conspiracies in the 2020 US presidential election.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+16.
+908–919.
+- Sharma et al.
+(2022b) Karishma Sharma, Yizhou Zhang, and Yan Liu.
+2022b.
+COVID-19 vaccine misinformation campaigns and social media narratives.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+16.
+920–931.
+- Solovev and Pröllochs (2022) Kirill Solovev and Nicolas Pröllochs.
+2022.
+Moral emotions shape the virality of COVID-19 misinformation on social media.
+In Proceedings of the ACM Web Conference 2022.
+3706–3717.
 - Suarez-Lledo and Alvarez-Galvez (2021) Victor Suarez-Lledo and Javier Alvarez-Galvez.
@@
 Journal of Medical Internet Research 23, 1 (2021), e17187.
+- Taira et al.
+(2021) Breena R Taira, Vanessa Kreger, Aristides Orue, and Lisa C Diamond.
+2021.
+A pragmatic assessment of Google Translate for emergency department instructions.
+Journal of General Internal Medicine 36, 11 (2021), 3361–3365.
 - Tang et al.
@@
 Gopal, and Guanglei Zhang.
-2023.
+2024.
 Racial discrimination and anti-discrimination: The COVID-19 pandemic’s impact on Chinese restaurants in North America.
-Information Systems Research (2023).
-- Teng et al. (2022) Xian Teng, Yu-Ru Lin, Wen-Ting Chung, Ang Li, and Adriana Kovashka. 2022. Characterizing user susceptibility to COVID-19 misinformation on Twitter. In ICWSM.
+Information Systems Research 35, 3 (2024), 1274–1295.
+- Teng et al.
+(2022) Xian Teng, Yu-Ru Lin, Wen-Ting Chung, Ang Li, and Adriana Kovashka.
+2022.
+Characterizing user susceptibility to COVID-19 misinformation on Twitter.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+16.
+1005–1016.
 - Van Prooijen et al.
@@
 - Vosoughi et al. (2018) Soroush Vosoughi, Deb Roy, and Sinan Aral. 2018. The spread of true and false news online. Science 359, 6380 (2018), 1146–1151.
-- Wang and Rzeszotarski (2023) Luping Wang and Jeffrey M. Rzeszotarski. 2023. Understanding motivational factors in social media news sharing decisions. In CSCW.
+- Wang and Rzeszotarski (2023) Luping Wang and Jeffrey M.
+Rzeszotarski.
+2023.
+Understanding motivational factors in social media news sharing decisions.
+Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (2023), 1–30.
 - Wang et al.
@@
 Journal of the Association for Information Science and Technology 73, 5 (2022), 726–737.
-- Wang et al. (2023a) Yuping Wang, Chen Ling, and Gianluca Stringhini. 2023a. Understanding the use of images to spread COVID-19 misinformation on Twitter. In CSCW.
 - Wang et al.
-(2023b) Ying Wang, Hongwei Zhang, Wang Gao, and Cai Yang.
-2023b.
+(2023) Ying Wang, Hongwei Zhang, Wang Gao, and Cai Yang.
+2023.
 Spillover effects from news to travel and leisure stocks during the COVID-19 pandemic: Evidence from the time and frequency domains.
 Tourism Economics 29, 2 (2023), 460–487.
-- Yan et al. (2022) Muheng Yan, Yu-Ru Lin, and Wen-Ting Chung. 2022. Are mutated misinformation more contagious? A case study of COVID-19 misinformation on Twitter. In WebSci.
+- Yan et al.
+(2022) Muheng Yan, Yu-Ru Lin, and Wen-Ting Chung.
+2022.
+Are mutated misinformation more contagious?
+A case study of COVID-19 misinformation on Twitter.
+In Proceedings of the 14th ACM Web Science Conference 2022.
+336–347.
+- Zannettou (2021) Savvas Zannettou.
+2021.
+”I Won the Election!”: An empirical analysis of soft moderation interventions on Twitter.
+In Proceedings of the International AAAI Conference on Web and Social Media, Vol.
+15.
+865–876.
+- Zavolokina et al.
+(2024) Liudmila Zavolokina, Kilian Sprenkamp, Zoya Katashinskaya, Daniel Gordon Jones, and Gerhard Schwabe.
+2024.
+Think fast, think slow, think critical: Designing an automated propaganda detection tool.
+In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems.
+1–24.
+- Zhou et al.
+(2021) Cheng Zhou, Kai Li, and Yanhong Lu.
+2021.
+Linguistic characteristics and the dissemination of misinformation in social media: The moderating effect of information richness.
+Information Processing & Management 58, 6 (2021), 102679.
 Supplementary Materials
+for “From News Source Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis”
+Appendix S1 Estimation Results for Main Analysis
+The full estimation results for the analysis in the main paper are reported in Tables S1–S3.
+Specifically, Table S1 reports the estimation results for the topic diversity.
+Table S2 reports the estimation results for the engagement at news source sharer level.
+Table S3 reports the estimation results for the engagement at post viewer level.
+Additionally, Table S4 reports the estimation results for the subset analysis of post viewer engagement within false news posts.
+| (1) | (2) | (3) | |
+| News claim | News post | Claim Post | |
+| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | |
+| Verified | ∗∗∗ | ||
+| AccountAge | |||
+| Followers | ∗∗∗ | ||
+| Followees | ∗∗∗ | ∗∗∗ | |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✗ | ✓ | ✓ |
+| 0.179 | ✗ | ✗ |
+| (1) | (2) | |
+| News lifetime | Post count | |
+| Falsehood | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗ | |
+| Diversity | ||
+| Falsehood Diversity | ||
+| Words | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ |
+| Language | ✓ | ✓ |
+| MonthYear | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ |
+| 0.481 | 0.188 |
+| (1) | (2) | (3) | |
+| Repost count | Like count | Reply count | |
+| Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
+| Anger | ∗∗∗ | ∗∗∗ | ∗∗ |
+| Disgust | ∗∗ | ||
+| Fear | ∗ | ∗ | |
+| Joy | ∗ | ∗∗∗ | |
+| Sadness | |||
+| Surprise | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✓ | ✓ | ✓ |
+| (1) | (2) | (3) | |
+| Repost count | Like count | Reply count | |
+| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Diversity | ∗∗∗ | ∗∗∗ | ∗∗ |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗ | |
+| News-level REs | ✓ | ✓ | ✓ |
 Appendix S2 Cross-Correlations & Variance Inflation Factors
-We additionally examine the cross-correlations and variance inflation factors among the independent variables to mitigate the concern about multicollinearity.
-Table S1 reports the cross-correlations among the variables in news items.
-Table S2 reports the cross-correlations among the variables in posts.
-The cross-correlations in in news items and posts are fairly small.
-Additionally, the variance inflation factors (VIFs) are reported in Table S3.
+We provide the cross-correlations to show how the independent variables relate to each other at the news claim and news post levels.
+Table S5 reports the cross-correlations among the variables in news claims.
+Table S6 reports the cross-correlations among the variables in news posts.
+Additionally, we examine the Variance Inflation Factors (VIFs) among the independent variables to mitigate the concern about multicollinearity.
+The variance inflation factors are reported in Table S7.
 We find that the variance inflation factors for the independent variables are all close to one and well below the critical threshold of four, which indicates that multicollinearity is not an issue of our analysis.
@@
 | Mean VIF |
+Appendix S3 Robustness Check With Original 26 Topics
+We adopt the original 26 topics and associated keywords directly from previous research (Chandrasekaran et al., 2020) and repeat the analysis.
+Table S8 reports the estimation results for the topic diversity.
+Table S9 reports the estimation results for the engagement at news source sharer level.
+Table S10 reports the estimation results for the engagement at post viewer level.
+All the results are robust and consistently support our main findings.
+| (1) | (2) | (3) | |
+| News claim | News post | Claim Post | |
+| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗ | ||
+| Verified | ∗∗∗ | ||
+| AccountAge | |||
+| Followers | ∗∗∗ | ||
+| Followees | ∗∗∗ | ∗∗∗ | |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✗ | ✓ | ✓ |
+| 0.196 | ✗ | ✗ |
+| (1) | (2) | |
+| News lifetime | Post count | |
+| Falsehood | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗ | |
+| Diversity | ||
+| Falsehood Diversity | ||
+| Words | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ |
+| Language | ✓ | ✓ |
+| MonthYear | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ |
+| 0.478 | 0.185 |
+| (1) | (2) | (3) | |
+| Repost count | Like count | Reply count | |
+| Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Falsehood Diversity | |||
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✓ | ✓ | ✓ |
 Appendix S4 Robustness Check at Theme Level
@@
 Table S11 reports the estimation results for the topic diversity.
-Table S5 reports the estimation results for the engagement at news sharer level.
+Table S12 reports the estimation results for the engagement at news source sharer level.
 Table S13 reports the estimation results for the engagement at post viewer level.
@@
 | (1) | (2) | (3) | |
-News | Post | News Post | |
+| News claim | News post | Claim Post | |
 | Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
@@
 | News-level REs | ✗ | ✓ | ✓ |
-0.147 | ✗ | ✗ |
+| 0.147 | ✗ | ✗ |
 | (1) | (2) | |
-Lifetime | Post Count | |
+| News lifetime | Post count | |
 | Falsehood | ∗∗∗ | ∗∗∗ |
@@
 | Intercept | ∗∗∗ | ∗∗∗ |
-0.480 | 0.188 |
+| 0.480 | 0.188 |
 | (1) | (2) | (3) | |
@@
 To reduce potential unobserved confounding factors and ensure the robustness of our results, we further conduct propensity score matching for true and false news claims.
-Specifically, we match true and false news claims based on the variables of emotions and the number of words, with the caliper of 0.3, “common support”, and “noreplacement” settings.
-Consequently, we discard 30 false news claims and achieve all biases lower than 0.05.
-Subsequently, we repeat our analysis based on the match true and false news claims and their associated posts on X.
+Specifically, we match true and false news claims based on the variables of emotions and the number of words, with settings including the caliper of 0.3, “common support”, and “noreplacement.” Consequently, we discard 30 false news claims and achieve all biases lower than 0.05.
+Subsequently, we repeat our analysis based on the matched true and false news claims and their associated posts on X.
 Table S14 reports the estimation results for the topic diversity.
-Table S8 reports the estimation results for the engagement at news sharer level.
+Table S15 reports the estimation results for the engagement at news source sharer level.
 Table S16 reports the estimation results for the engagement at post viewer level.
@@
 | (1) | (2) | (3) | |
-News | Post | News Post | |
+| News claim | News post | Claim Post | |
 | Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
@@
 | News-level REs | ✗ | ✓ | ✓ |
-0.181 | ✗ | ✗ |
+| 0.181 | ✗ | ✗ |
 | (1) | (2) | |
-Lifetime | Post Count | |
+| News lifetime | Post count | |
 | Falsehood | ∗∗∗ | ∗∗∗ |
@@
 | Intercept | ∗∗∗ | ∗∗∗ |
-0.482 | 0.188 |
+| 0.482 | 0.188 |
 | (1) | (2) | (3) | |
@@
 | News-level REs | ✓ | ✓ | ✓ |
-Appendix S4 Robustness Check with BERTopic
-To ensure the generalization of our results across the topic modeling methods, we use BERTopic, a popular unsupervised topic modeling tool, as an alternative method to automatically cluster topics.333https://maartengr.github.io/BERTopic/index.html The BERTopic model generates topic clusters for the whole news claims and posts.
-We use the topic clusters and their corresponding probabilities to calculate topic diversity again and repeat the analysis.
+Appendix S6 Robustness Check With Inclusion of QAnon Conspiracy Theory
+We further include QAnon to extend conspiracy theories and repeat our analysis.
+QAnon is a far-right conspiracy theory alleging that a cabal of politicians and media personalities are engaged in a war to undermine society (Papasavva et al., 2022; Bär et al., 2023; Kim and Kim, 2023).
+The QAnon discourse has become more diverse and evolved into a global phenomenon during the COVID-19 pandemic (Hoseini et al., 2023).
+We adopt a list of QAnon-specific keywords from previous research (Sharma et al., 2022a) to identify QAnon conspiracy theory in our dataset.
+The performance of this keyword list in identifying QAnon discourse has been manually validated, showing high agreement with human assessments (Bär et al., 2023).
+As a result, we identify 6 QAnon news claims and 302 QAnon news posts.
 Table S17 reports the estimation results for the topic diversity.
-Table S11 reports the estimation results for the engagement at news sharer level.
+Table S18 reports the estimation results for the engagement at news source sharer level.
 Table S19 reports the estimation results for the engagement at post viewer level.
@@
 | (1) | (2) | (3) | |
-News | Post | News Post | |
+| News claim | News post | Claim Post | |
+| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | |
+| Verified | ∗∗∗ | ||
+| AccountAge | |||
+| Followers | ∗∗∗ | ||
+| Followees | ∗∗∗ | ∗∗∗ | |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✗ | ✓ | ✓ |
+| 0.179 | ✗ | ✗ |
+| (1) | (2) | |
+| News lifetime | Post count | |
+| Falsehood | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗∗ | |
+| Diversity | ||
+| Falsehood Diversity | ||
+| Words | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ |
+| Language | ✓ | ✓ |
+| MonthYear | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ |
+| 0.481 | 0.188 |
+| (1) | (2) | (3) | |
+| Repost count | Like count | Reply count | |
+| Falsehood | ∗∗∗ | ∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Falsehood Diversity | ∗∗∗ | ∗∗ | ∗∗ |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✓ | ✓ | ✓ |
+Appendix S7 Robustness Check With Exclusion of Potential Debunking Posts
+News posts can also debunk the linked news items.
+We use keywords of “misinformation,” “misleading,” “fake,” “false,” and “disinformation” to identify potential debunking posts, resulting in news posts.
+We exclude these potential debunking posts and repeat our analysis.
+Table S20 reports the estimation results for the topic diversity.
+Table S21 reports the estimation results for the engagement at news source sharer level.
+Table S22 reports the estimation results for the engagement at post viewer level.
+All the results are robust and consistently support our main findings.
+| (1) | (2) | (3) | |
+| News claim | News post | Claim Post | |
+| Falsehood | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗∗∗ | ∗∗∗ | |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗ | ∗∗∗ | |
+| Verified | ∗∗∗ | ||
+| AccountAge | |||
+| Followers | ∗∗∗ | ||
+| Followees | ∗∗∗ | ∗∗ | |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✗ | ✓ | ✓ |
+| 0.173 | ✗ | ✗ |
+| (1) | (2) | |
+| News lifetime | Post count | |
+| Falsehood | ∗∗∗ | ∗∗∗ |
+| Conspiracy | ∗ | |
+| Diversity | ||
+| Falsehood Diversity | ||
+| Words | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ |
+| Language | ✓ | ✓ |
+| MonthYear | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ |
+| 0.496 | 0.191 |
+| (1) | (2) | (3) | |
+| Repost count | Like count | Reply count | |
+| Falsehood | ∗∗ | ∗∗∗ | |
+| Conspiracy | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Diversity | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Falsehood Diversity | ∗∗∗ | ∗ | ∗∗ |
+| Words | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Media | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Verified | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| AccountAge | ∗∗∗ | ∗∗∗ | ∗∗ |
+| Followers | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Followees | ∗∗∗ | ∗∗∗ | ∗∗∗ |
+| Emotions | ✓ | ✓ | ✓ |
+| Language | ✓ | ✓ | ✓ |
+| MonthYear | ✓ | ✓ | ✓ |
+| Intercept | ∗∗∗ | ∗∗∗ | |
+| News-level REs | ✓ | ✓ | ✓ |
+Appendix S8 Robustness Check With BERTopic
+To ensure the generalization of our results across the topic modeling methods, we use BERTopic, a popular unsupervised topic modeling tool, as an alternative method to automatically cluster topics.999https://maartengr.github.io/BERTopic/index.html The BERTopic model generates topic clusters for the whole news claims and posts.
+We use the topic clusters and their corresponding probabilities to calculate topic diversity again and repeat the analysis.
+Table S23 reports the estimation results for the topic diversity.
+Table S24 reports the estimation results for the engagement at news source sharer level.
+Table S25 reports the estimation results for the engagement at post viewer level.
+All the results are robust and consistently support our main findings.
+| (1) | (2) | (3) | |
+| News claim | News post | Claim Post | |
 | Falsehood | ∗∗ | ||
@@
 | News-level REs | ✗ | ✓ | ✓ |
-0.041 | ✗ | ✗ |
+| 0.041 | ✗ | ✗ |
 | (1) | (2) | |
-Lifetime | Post Count | |
+| News lifetime | Post count | |
 | Falsehood | ∗∗∗ | ∗∗∗ |
@@
 | Intercept | ∗∗∗ | ∗∗∗ |
-0.508 | 0.192 |
+| 0.508 | 0.192 |
 | (1) | (2) | (3) | |
@@
 | News-level REs | ✓ | ✓ | ✓ |
+Appendix S9 Additional Analysis on Quote Count
+Table S26 reports the full estimation results for quote count.
+| Quote count | |
+| Falsehood | |
+| Conspiracy | ∗∗∗ |
+| Diversity | ∗∗∗ |
+| Falsehood Diversity | ∗∗∗ |
+| Words | ∗∗∗ |
+| Media | ∗∗∗ |
+| Verified | ∗∗∗ |
+| AccountAge | ∗∗∗ |
+| Followers | ∗∗∗ |
+| Followees | ∗∗∗ |
+| Emotions | ✓ |
+| Language | ✓ |
+| MonthYear | ✓ |
+| Intercept | ∗∗∗ |
+| News-level REs | ✓ |